### v1

# A polynomial bound on the pathwidth of graphs edge-coverable by $k$ shortest paths

Abstract
Dumas, Foucaud, Perez and Todinca (2024) recently proved that every graph whose edges can be covered by $k$ shortest paths has pathwidth at most $O(3^{k})$. In this paper, we improve this upper bound on the pathwidth to a polynomial one; namely, we show that every graph whose edge set can be covered by $k$ shortest paths has pathwidth $O(k^{4})$, answering a question from the same paper. Moreover, we prove that when $k\leq 3$, every such graph has pathwidth at most $k$ (and this bound is tight). Finally, we show that even though there exist graphs with arbitrarily large treewidth whose vertex set can be covered by $2$ isometric trees, every graph whose set of edges can be covered by $2$ isometric trees has treewidth at most $2$.
1 Introduction
Graph covering is an old and recurrent topic in graph theory. For a fixed class $\mathcal{H}$ of graphs, the covering literature aims to answer two types of problems. First are optimization problems: given a graph $G$, what is the minimum number of graphs from $\mathcal{H}$ needed to “cover” [^1] Various notions of coverings can usually be considered here. $G$? We refer to [KU16] for a more general overview on the topic. The second type are structural problems: for the fixed class $\mathcal{H}$ of graphs and some $k\geq 1$, what is the structure of the graphs that can be “covered” by (at most) $k$ graphs from $\mathcal{H}$?
We say that a graph $G$ is edge-coverable (resp. vertex-coverable) by some subgraphs $H_{1},\ldots,H_{k}$ if every edge (resp. vertex) of $G$ belongs to at least one of the subgraphs $H_{i}$. Motivated by some algorithmic applications to the Isometric Path Cover problem and its variants, Dumas, Foucaud, Perez and Todinca [DFPT24] recently worked on a particular metric version of such covering problems. They proved that every graph edge-coverable (resp. vertex-coverable) by few shortest paths shares structural similarities with a simple path. More precisely, if a graph is edge-coverable or vertex-coverable by $k$ shortest paths, its pathwidth is smaller than an exponential in $k$ (see Section˜2 for a definition of pathwidth).
Theorem 1.1 ([DFPT24]).
Let $k\geq 1$ and $G$ be a graph vertex-coverable by $k$ shortest paths. Then $\mathrm{pw}(G)=O(k\cdot 3^{k})$. Moreover, if $G$ is edge-coverable by $k$ shortest paths, then $\mathrm{pw}(G)=O(3^{k})$.
The proof of Theorem˜1.1 uses a branching approach to show that in every graph vertex-coverable (resp. edge-coverable) by at most $k$ shortest paths, for every vertex $u$ and every $i\geq 1$, the number of vertices at distance exactly $i$ from $u$ in $G$ is $O(k\cdot 3^{k})$ (resp. $O(3^{k})$). We also mention that a “coarse” variant of Theorem˜1.1 has recently been proved in [HP25]. Then, the bags of the path-decomposition proving Theorem˜1.1 are the union of every two consecutive layers in a BFS-layering of $G$ (starting from an arbitrary vertex). Such a path-decomposition is clearly not optimal in general: if we let $S_{h}$ denote the star with $h$ leaves, then $\mathrm{pw}(S_{h})=1$, while such a decomposition will always have width at least $\lfloor\tfrac{h}{2}\rfloor$, whatever the choice of $u$ is. The authors of [DFPT24] asked if the bounds from Theorem˜1.1 can be made polynomial in $k$. Our main result is a positive answer in the edge-coverable case.
Theorem 1.2.
Let $k\geq 1$, and $G$ be a graph edge-coverable by $k$ shortest paths. Then $\mathrm{pw}(G)=O(k^{4})$.
Our proof uses a different approach from [DFPT24]. Roughly speaking, it consists in finding some separator $X$ of polynomial size, such that for every component $C$ of GXG-X, there exists one path $P$ from the covering set of paths, such that $C$ has a “nice” layering with respect to $P$.
In terms of lower bounds, there exist for arbitrary values of $k\geq 1$ some graphs edge-coverable by $k$ shortest paths and with pathwidth equal to $k$ [DFPT24]. However, note that no better lower bound is known. Our second result is that this lower bound is tight for $k\leq 3$.
Theorem 1.3.
Let $k\leq 3$, and $G$ be a graph edge-coverable by $k$ shortest paths. Then $\mathrm{pw}(G)\leq k$.
Given a graph $G$, a subgraph $H$ of $G$ is isometric if every shortest path in $H$ is also a shortest path in $G$. For instance, shortest paths in $G$ are isometric subgraphs of $G$.
In view of the previous results, the following question naturally arises: given a family $\mathcal{H}$ of graphs, what is the structure of a graph $G$ edge/vertex-coverable by at most $k$ isometric subgraphs all isomorphic to graphs in $\mathcal{H}$? For instance, if $\mathcal{H}$ is the family of all trees, does $G$ somehow look like a tree?
Extending a result of Aigner and Fromme [AF84] for paths, Ball, Bell, Guzman, Hanson-Colvin and Schonsheck [BBG+17] proved that every graph vertex-coverable by $k$ isometric subtrees has cop number at most $k$. In particular, it is well-known that the cop number of a graph is always upper-bounded by its treewidth (see Section˜5 for a definition of treewidth). In analogy to Theorem˜1.1, a natural question is then the following: is it true that every graph which is vertex/edge-coverable by a small number of isometric subtrees has small treewidth? This question turns out to have a negative answer when considering vertex-coverability, as shown by the following example, pointed to us by Marcin Briański (personal communication): for every $t\geq 1$, let $G_{t}$ be the graph on $2t+2$ vertices obtained from the biclique $K_{t,t}$ after adding two vertices $a,b$ of degree $t$ respectively adjacent to all vertices from each of the sides of $K_{t,t}$ (see Figure˜1 for the case $t=3$). Then, $G_{t}$ is vertex-coverable by the two induced stars $S_{a},S_{b}$ respectively centered in $a$ and $b$ (which are in particular isometric subgraphs of $G_{t}$), but has treewidth at least $t$, as it contains $K_{t,t}$ as a minor. Our third result implies that such a construction does not exist anymore when considering graphs which are edge-coverable by two isometric trees.
Figure 1: This graph is vertex-coverable by $2$ isometric trees (one with red vertices, the other with blue vertices) and contains $K_{3,3}$ as a minor (black edges). Hence, it has treewidth at least $3$.
Theorem 1.4.
Let $G$ be a graph which is edge-coverable by $2$ isometric trees. Then $\mathrm{tw}(G)\leq 2$.
Algorithmic motivations.
Routing problems play a central role not only in algorithmic graph theory and complexity, but also in structural graph theory. One of the best examples is the Disjoint Paths problem, appearing in the seminal Graph Minor series of papers of Robertson and Seymour [RS95], in which one is given as an input a graph $G$, an integer $k$ and $k$ pairs of vertices, and one must decide whether there exists $k$ pairwise disjoint paths connecting all pairs of vertices. This problem has been shown to be solvable in FPT time, when parameterized by the number $k$ of shortest paths (see [RS95, KKR12]), and this result has been a key ingredient to design efficient algorithms deciding the existence of minors in graphs. The metric variant Disjoint Shortest Path of this problem introduced in [ET98] (in which one further wants to decide whether there exist $k$ pairwise disjoint shortest paths connecting all input pairs of vertices) has attired a lot of attention, and was recently proved to be solvable in XP time, when parameterized by $k$ [BK17, Loc21, BNRZ23]. [^2] It is worth mentioning that all the aforementioned problems are NP-hard [FHW80, Loc21].
A number of metric variants of routing problems of similar flavour has been studied over the last decades. Among them, Isometric Path Cover studied in [AF84, FF01] takes as input a graph $G$ and an integer $k$, and asks whether $G$ is vertex-coverable by $k$-shortest paths. Its routing version Isometric Path Cover With Terminals takes as input a graph $G$, some integer $k$, and $k$ pairs of vertices, and asks whether there exist $k$ shortest paths that vertex-cover $G$ and connect all the input pairs of vertices. Both problems are NP-Complete [CDD+22, DFPT24]. Combining Theorem˜1.1 with Courcelle’s theorem [Cou90], the authors from [DFPT24] proved the existence of some algorithms running in XP and FPT time when parameterized by $k$, to solve respectively Isometric Path Cover and Isometric Path Cover With Terminals. In particular, it is still open whether Isometric Path Cover problem admits an FPT algorithm parameterized by $k$ [DFPT24]. We also refer to [CCFV26] for more results on Isometric Path Cover, and to [CDF+25] for some recent work on similar metric versions of routing problems.
A number of metric variants of routing problems of similar flavour has been studied over the last decades. Among them, Isometric Path Cover studied in [AF84, FF01] takes as input a graph $G$ and an integer $k$, and asks whether $G$ is vertex-coverable by $k$-shortest paths. Its routing version Isometric Path Cover With Terminals takes as input a graph $G$, some integer $k$, and $k$ pairs of vertices, and asks whether there exist $k$ shortest paths that vertex-cover $G$ and connect all the input pairs of vertices. Both problems are NP-Complete [CDD+22, DFPT24]. Combining Theorem˜1.1 with Courcelle’s theorem [Cou90], the authors from [DFPT24] proved respectively the existence of some algorithms running in XP and FPT time when parameterized by $k$, to solve respectively Isometric Path Cover and Isometric Path Cover With Terminals. In particular, it is still open whether Isometric Path Cover problem admits an FPT algorithm parameterized by $k$ [DFPT24]. We also refer to [CDF+25] for some recent work on similar metric versions of routing problems. It is worth mentioning that Isometric Path Cover has be shown to be solvable in FPT time when parameterized by the isometric path complexity of a graph [CCFV26], a recent metric parameter of graphs introduced in [CDD+22] (under the name of isometric path antichain cover number).
As mentioned above, the original parameterized algorithms from [DFPT24] solving Isometric Path Cover and its variant with terminals rely on Courcelle’s theorem, and thus the dependencies in $k$ are highly non-trivial and do not give rise in general to practical algorithms. In [CDF+25, Section 3], the authors designed a dynamic programming procedure solving the partition variant Isometric Path Partition of Isometric Path Cover in time $n^{O(\mathrm{tw}(G))}$, where $\mathrm{tw}(G)$ denotes the treewidth of the input graph. As pointed to us by Florent Foucaud (personal communication), it is very likely that combining our bound from Theorem˜1.2 with such an approach could lead to better running times than the ones given by Courcelle’s theorem for algorithms solving Isometric Path Edge-Cover and its version with terminals (the natural edge-covering variant of Isometric Path Cover).
Organization of the paper.
Section˜2 contains all preliminary definitions and some basic results related to pathwidth. Section˜3 contains a proof of Theorem˜1.2, Section˜4 contains a proof of Theorem˜1.3, and Section˜5 contains a proof of Theorem˜1.4. Note that Sections˜3, 4 and 5 are completely independent from each other. Section˜6 contains some discussion and additional questions left open by this work.
2 Preliminaries
2.1 Notations and Basic Definitions
Graphs.
In this work, we consider undirected finite simple graphs, without loops. For a graph $G$, we denote with $V(G)$ its set of vertices, and $E(G)$ its set of edges. For any set $X\subseteq V(G)$ of vertices, $G[X]$ denotes the subgraph of $G$ induced by $X$, that is, the vertex set of $G[X]$ is $X$, and $G[X]$ contains all edges of $G$ with both end-points in $X$. GXG-X denotes the graph induced by $G$ on $V(G)\setminus X$. For simplicity, for every vertex $x$, we let GxG-x denote the graph $G-\{x\}$. For every subsets $X,Y\subseteq V(G)$ of vertices, we denote $E_{G}[X,Y]:=\{xy\in E(G):x\in X\setminus Y,y\in Y\setminus X\}$. Let $G^{\prime}$ be a graph, whose vertex set potentially intersects the one of $G$. The union of $G$ and $G^{\prime}$ is the graph $G\cup G^{\prime}:=(V(G)\cup V(G^{\prime}),E(G)\cup E(G^{\prime}))$.
For every $v\in V(G)$, we let $N_{G}(v)$ (or simply $N(v)$ when the context is clear) denote the set of neighbors of $v$ in $G$. Similarly, for every subset $X\subseteq V(G)$ of vertices, we let $N_{G}(X)$ (or simply $N(X)$) denote the set of vertices in $V(G)\setminus X$ having at least one neighbor in $X$. If $U\subseteq V(G)$, we similarly let $N_{U}(X)$ denote the set of vertices in $U\setminus X$ having at least one neighbor in $X$. For every $v\in V(G)$, we let $N_{G}(v)$ (or simply $N(v)$ when the context is clear) denote the set of neighbors of $v$. Similarly, for every subset $X\subseteq V(G)$ of vertices, we let $N_{G}(X)$ (or simply $N(X)$) denote the set of vertices in $V(G)\setminus X$ having at least one neighbor in $X$. If $U\subseteq V(G)$, we similarly let $N_{U}(X)$ denote the set of vertices in $U\setminus X$ having at least one neighbor in $X$.
For every three subsets $X,A,B\subseteq V(G)$, we say that $X$ separates $A$ from $B$ in $G$ if no component of GXG-X contains both a vertex from $A$ and $B$.
Paths.
A path $P$ is a sequence of pairwise distinct vertices $v_{1},\ldots,v_{k}$ such that for each $i<k$, $v_{i}v_{i+1}\in E(G)$. A path connecting two vertices $v_{1}=x,v_{k}=y$ will also be called an $xy$-path. We call the vertices $v_{1},v_{k}$ the endvertices of $P$, while all its other vertices are its internal vertices. The length of $P$ corresponds to its number of edges. For each $i\leq j$, we let $P[v_{i},v_{j}]$ denote the subpath $v_{i},v_{i+1},\ldots,v_{j}$ of $P$, and we let $P^{-1}$ denote the path $v_{k},\ldots,v_{1}$. For every two paths $P=v_{1},\ldots,v_{p}$ and $Q=w_{1},\ldots,w_{q}$ such that $v_{p}=w_{1}$, and such that $P-v_{p}$ is disjoint from $Q-w_{1}$, we define the path $P\cdot Q:=v_{1},\ldots,v_{p}=w_{1},\ldots,w_{q}$. If $P=v_{1},\ldots,v_{k}$ is a path, for every $1\leq i<j\leq k$, we say that $v_{i}$ is before $v_{j}$ on $P$, and that $v_{j}$ is after $v_{i}$ on $P$. Note that by definition, a path is a sequence of vertices; however, by abuse of notations, we will often amalgamate a path $P=v_{1},\ldots,v_{k}$ of $G$ with its corresponding subgraph in $G$ (that is, the graph with vertices $v_{1},\ldots,v_{k}$ and edges $v_{i}v_{i+1}$ for each $i$). A path $P$ is vertical with respect to a vertex $u$, if for every $i\in\mathbb{N}$, $P$ has at most one vertex at distance $i$ from $u$ in $G$. Note that every path vertical with respect to some vertex $u$ is a shortest path, and that every shortest path starting from $u$ is vertical with respect to $u$.
Covers.
Let $G$ be a graph and $H_{1},\ldots,H_{k}$ be subgraphs of $G$. We say that the graphs $H_{1},\ldots,H_{k}$ edge-cover $G$ if $E(G)=\bigcup_{i=1}^{k}E(H_{i})$. Similarly, we say that $H_{1},\ldots,H_{k}$ vertex-cover $G$ if $V(G)=\bigcup_{i=1}^{k}V(H_{i})$.
Metrics.
We let $d_{G}(\cdot,\cdot)$ (or simply $d(\cdot,\cdot)$ when the context is clear) denote the shortest-path metric of a graph $G$. Recall that a graph $H$ is an isometric subgraph of $G$ if $H$ is a subgraph of $G$ such that every shortest path in $H$ is also a shortest path in $G$. For every vertex $v$ of a graph $G$, we let $\mathrm{ecc}(v):=\max\{d_{G}(u,v):u\in V(G)\}\in\mathbb{N}\cup\{\infty\}$ denote the eccentricity of $v$.
Orders.
Let $(X,\leq)$ be a partially ordered set. For every two disjoint subsets $A,B\subseteq X$, we write $A<B$ if $a<b$ for every $(a,b)\in A\times B$. Moreover, for every $x\in X$ and every $A\subseteq X$, we let $A|_{\leq x}:=\{a\in A:a\leq x\}$ and $A|_{\geq x}:=\{a\in A:a\geq x\}$. We define similarly $A|_{<x}$ and $A|_{>x}$. A chain is a sequence $(x_{1},\dots,x_{k})$ of distinct elements of $X$ such that $x_{1}\leq x_{2}\leq\dots\leq x_{k}$.
2.2 Path-decompositions
We collect in this section a number of basic properties of pathwidth.
Path-decomposition.
A path-decomposition of a graph $G$ is a sequence $\mathcal{P}=(X_{1},X_{2},\ldots,X_{q})$ of vertex subsets of $G$, called bags, such that for every edge $\{x,y\}\in E$ there is at least one bag containing both endpoints, and for every vertex $x\in V$, the bags containing $x$ form a continuous subsequence of $\mathcal{P}$. The width of $\mathcal{P}$ is $\max\{\lvert X_{i}\rvert-1:1\leq i\leq q\}$, and the pathwidth $\mathrm{pw}(G)$ of $G$ is the smallest width of a decomposition, among all path-decomposition of $G$. The bags $X_{1},X_{q}$ are called the extremities of $\mathcal{P}$. We say that $X_{1}$ is the leftmost bag, and that $X_{q}$ is the rightmost bag.
Remark 2.1.
For every graph $G$ and $X\subseteq V(G)$, $\mathrm{pw}(G)\leq\lvert X\rvert+\mathrm{pw}(G-X)$.
Vertex-separation number.
Let $G$ be a graph and << denote a total ordering on $V(G)$. The vertex-separation number of $(G,<)$ is the value $\min_{w\in V(G)}|\{u\in V(G):u<w~\text{and}~\exists v\geq w,uv\in E(G)\}|$. The vertex separation number of $G$ is the minimum over the vertex separation numbers of $(G,<)$, for all possible linear orderings << of $G$.
Theorem 2.2 ([Kin92]).
The vertex separation number of a graph equals its pathwidth.
Layerings.
A layering of a graph $G$ is a partition of its vertices into a sequence $(V_{1},\ldots,V_{t})$ of subsets of $V(G)$ called layers, such that each graph $G[V_{i}]$ is an independent set, and for every $i,j\in\{1,\ldots,t\}$ such that $|j-i|\geq 2$, we have $E_{G}(V_{i},V_{j})=\emptyset$. [^3] Note that our definition of layering differs from the one which is commonly used in the literature, as one usually allows in a layering edges between vertices from a same layer.
A $k$-layering of $G$ is a layering $(V_{1},\ldots,V_{t})$ such that for every $i<t$, we have $\lvert E_{G}(V_{i},V_{i+1})\rvert\leq k$. Note that in particular, it implies that $\lvert N_{V_{i+1}}(V_{i})\rvert\leq k$ and $\lvert N_{V_{i}}(V_{i+1})\rvert\leq k$. The bags $V_{1}$ and $V_{t}$ are the extremities of the layering.
Lemma 2.3.
Let $G$ be a graph admitting a $k$-layering. Then $G$ has pathwidth at most $k$. Moreover, if $G$ is connected, then $G$ has a path-decomposition of width at most $k$, whose extremities are exactly the extremities of the $k$-layering.
Proof.
We let $(V_{1},\ldots,V_{t})$ denote a $k$-layering of $G$, and assume without loss of generality that $G$ has no isolated vertex. We also set $V_{0}:=V_{t+1}:=\emptyset$. For every $i\in\{1,\ldots,t\}$, we set $V_{i}^{\ell}:=N_{V_{i}}(V_{i-1})$ and $V_{i}^{r}:=N_{V_{i}}(V_{i+1})$. Note in particular that, as $G$ has no isolated vertex, for each $i\in\{1,\ldots,t\}$ we have $V_{i}=V_{i}^{\ell}\cup V_{i}^{r}$, $|V_{i}^{\ell}|\leq k$, $|V_{i}^{r}|\leq k$, and $V_{1}^{\ell}=V_{t}^{r}=\emptyset$, and that for each $i<t$, all the edges of $G$ between $V_{i}$ and $V_{i+1}$ are between $V_{i}^{r}$ and $V_{i+1}^{\ell}$.
We will define a total order $\leq$ on $V(G)$ such that $(G,\leq)$ has vertex separation number at most $k$. In particular, Theorem˜2.2 will then immediately imply that $\mathrm{pw}(G)\leq k$. We let $\leq$ be any total ordering on $V(G)$ satisfying the following two properties:
  • for every $i<j$, we have $V_{i}<V_{j}$;
  • for every $i\in\{1,\ldots,t\}$, we have $V^{\ell}_{i}\setminus V^{r}_{i}<V^{\ell}_{i}\cap V^{r}_{i}<V^{r}_{i}\setminus V^{\ell}_{i}$.
We now show that $(G,\leq)$ has vertex separation number at most $k$ and let $w\in V(G)$. We set $S_{w}:=\{u:u<w\text{ and }\exists v\geq w:uv\in E(G)\}$. Our goal is to show that $|S_{w}|\leq k$. We let $i$ denote the unique element from $\{1,\ldots,t\}$ such that $w\in V_{i}$. We define the sets $A:=V^{r}_{i}|_{<w},B:=V^{\ell}_{i}|_{\geq w}$ and $C:=V^{r}_{i-1}$. Note that $A,B,C$ are pairwise disjoint, and that moreover, as $(V_{1},\ldots,V_{t})$ is a layering of $G$, $A\subseteq S_{w}\subseteq A\uplus C$.
First, observe that if $B$ is empty, then we must have $S_{w}\cap C=\emptyset$, so that $S_{w}=A$, and thus as every vertex of $A$ is by definition of $V^{r}_{i}$ adjacent to a vertex of $V_{i+1}$, and as $|E_{G}(V_{i},V_{i+1})|\leq k$, we then have $|S_{w}|\leq k$. Assume now that $B$ is not empty. Then there exists an element $v\geq w$ such that $v\in V^{\ell}_{i}$. Thus, by definition of <<, we have $V_{i}|_{\leq w}\subseteq V^{\ell}_{i}$, and thus $A\subseteq V^{\ell}_{i}$.
We now set $k_{1}:=|E_{G}(C,A)|$ and $k_{2}:=|E_{G}(C,B)|$. As $A\subseteq V^{\ell}_{i}$, every element of $A$ must be incident to an edge from $E_{G}(A,C)$, hence $|A|\leq k_{1}$. As $C=V_{i-1}^{r}$, we have $|C|\leq k_{2}$. Moreover, note as $(V_{1},\ldots,V_{t})$ is a $k$-layering and as $A$ and $B$ are disjoint, we have $k_{1}+k_{2}\leq k$. In particular, it implies that
$|S_{w}|\leq|A|+|C|\leq k_{1}+k_{2}\leq k,$
hence we can conclude that $(G,\leq)$ has indeed vertex separation number at most $k$.
To prove the “Moreover” part of the lemma, observe that if $G$ is connected, then both $V_{1}$ and $V_{t}$ have size at most $k$. We claim that it is not hard to check that the path-decomposition given by Theorem˜2.2 can then be slightly modified so that its extremities are $V_{1}$ and $V_{k}$. ∎
3 A General Polynomial Bound
In this section, we prove Theorem˜1.2, which we restate here.
See 1.2
Throughout this section, we let $G$ be a graph whose edge set can be covered by a set $\mathcal{P}:=\{P_{1},\ldots,P_{k}\}$ of $k$ isometric paths. Our main result is the following, which will immediately imply a proof of Theorem˜1.2.
Theorem 3.1.
For every $i\in\{1,\ldots,k\}$, there exists a set of vertices $X_{i}\subseteq V$ of size at most $720k^{3}+4k$ such that every connected component of $G-X_{i}$ that intersects $P_{i}$ has pathwidth at most $6k$.
Proof that Theorem˜3.1 implies Theorem˜1.2.
For every $1\leq i\leq k$, we use Theorem˜3.1 to obtain the set $X_{i}$. We consider their union $X:=\bigcup_{i=1}^{k}X_{i}$. Observe that $\lvert X\rvert\leq\sum_{i\in[k]}\lvert X_{i}\rvert\leq k\cdot(720k^{3}+4k)=720k^{4}+4k^{2}$. We now let $C$ be a connected component of GXG-X. As $P_{1},\ldots,P_{k}$ edge-cover (and thus also vertex-cover) $G$, there is some $i\in\{1,\ldots,k\}$ such that the path $P_{i}$ intersects $C$. In particular, as $C$ is included in a connected component of $G-X_{i}$, it implies that $C$ has pathwidth smaller than $6k$, and thus that $\mathrm{pw}(G-X)\leq 6k$. Using Remark˜2.1 we conclude that $\mathrm{pw}(G)\leq\lvert X\rvert+\mathrm{pw}(G-X)\leq 720k^{4}+4k^{2}+6k$. ∎
The remainder of the section will consist in a proof of Theorem˜3.1. In Subsection 3.1 we prove some preliminary results, and show that our proof reduces to the problem of finding some special kind of separations in a subgraph $G_{0}$ of $G$ which is $6k$-layered (see Lemma˜3.9). In Subsection 3.2, we give a proof of Lemma˜3.9, which is the most technical part of our proof.
Without loss of generality, we will assume from now on that $i=1$, i.e., we will construct a set $X$ of size at most $720k^{3}+4k$ such that every component of GXG-X that intersects $P_{1}$ has pathwidth at most $6k$.
3.1 Preliminary Results
Parallel paths.
Let $P$ be a shortest path between two vertices $a,b$ of $G$. A path $Q$ of $G$ is parallel to $P$ in $G$ if $Q$ is a subpath of a shortest $ab$-path of $G$.
Note that this does not define a symmetric relation in general: the fact that a path $Q$ is parallel to another path $P$ does not necessarily imply that $P$ is parallel to $Q$. However, in the remainder of the proof, we will always consider the property of being parallel to the path $P_{1}$, which is fixed and connects vertices $a_{1}$ and $b_{1}$. In particular, the following simple observation implies that every graph which is edge-coverable by a few number of paths that are all parallel to $P_{1}$ has a very restrained structure.
Lemma 3.2.
Let $\{Q_{1},\ldots,Q_{\ell}\}$ be a collection of paths in $G$ that are all parallel to $P_{1}$. Then the graph $\bigcup_{j=1}^{\ell}Q_{j}$ is $\ell$-layered.
Proof.
We set $U:=\bigcup_{i=1}^{\ell}V(Q_{i})$, $r:=d(a_{1},b_{1})$, and for every $i\in\{0,\ldots,r\}$, we let $U_{i}:=\{u\in U:d(a_{1},u)=i\}$. As every path $Q_{j}$ is parallel to $P_{1}$, observe that for each $(i,j)\in\{0,\ldots,r\}\times\{1,\ldots,\ell\}$, we must have $|V(Q_{j})\cap U_{i}|\leq 1$. In particular, for each $i\in\{0,\ldots,r\}$, we have $|U_{i}|\leq\ell$. Hence $(U_{0},\ldots,U_{r})$ forms a partition of $U$, such that each edge of $G[U]$ has its two endvertices in some pair $(U_{i},U_{i^{\prime}})$ such that $|i^{\prime}-i|\leq 1$. Moreover, as the paths $Q_{j}$ are all parallel to $P_{1}$, no edge of a path $Q_{j}$ can have both endvertices in the same set $U_{i}$, so $(U_{0},\ldots,U_{r})$ is indeed a $\ell$-layering of $\bigcup_{j=1}^{\ell}Q_{j}$. ∎
Lemma 3.3.
If $P$ is a path parallel to $P_{1}$ with endvertices $u,v\in V(G)$, then every shortest $uv$-path in $G$ is also parallel to $P_{1}$.
Proof.
This immediately follows from the observation that every subpath of a shortest path is also a shortest path. ∎
We consider the partial order $\leq_{1}$ on $V(G)$ defined by setting for every two vertices $u,v\in V(G)$, $u\leq_{1}v$ if and only if $d_{G}(a_{1},u)\leq d_{G}(a_{1},v)$. Note in particular that the vertex set of every path $P$ which is parallel to $P_{1}$, forms a chain with respect to $\leq_{1}$.
Lemma 3.4.
Let $a\leq_{1}b\leq_{1}c$ be three vertices such that there exists two paths $P,Q$ both parallel to $P_{1}$, such that $P$ connects $a$ to $b$ and $Q$ connects $b$ to $c$. Then $P\cdot Q$ is parallel to $P_{1}$.
Proof.
Note that $P\cdot Q$ must be a shortest $ac$ path, as for each $i\geq 0$, it contains at most one vertex at distance exactly $i$ from $a_{1}$. In particular, it is not hard to conclude that $P\cdot Q$ is parallel to $P_{1}$. ∎
Reducing $\mathcal{P}$.
The first step in the proof of Theorem˜3.1 consists in modifying the covering family $\mathcal{P}$ in order to make it reduced (see definition below). The intuition behind this step is that our proof of Theorem˜3.1 works particularly well when all paths in $\mathcal{P}$ intersect $P_{1}$. Informally, the reduction operation consists in making $\mathcal{P}$ contain as much paths that intersect $P_{1}$ as possible. This step will turn out to be useful at the very end of our proof (see the proof of Lemma˜3.11).
A shortest path $Q$ of $G$ is called a reducing path if:
  • $Q$ is disjoint from $P_{1}$, and admits a subpath $P$ parallel to $P_{1}$, that connects two vertices $u,v$, such that $u<_{1}v$,
  • there exists a shortest path $R$ in $G$ from $u$ to $v$ that intersects $P_{1}$.
Figure˜2 depicts a reducing path. If a family $\mathcal{Q}$ of shortest paths has no reducing path, then we say that $\mathcal{Q}$ is reduced.
Proposition 3.5.
Let $\mathcal{Q}$ be a family of shortest paths. Then there exists a family $\mathcal{Q}^{\prime}$ of shortest paths which is reduced, such that $|\mathcal{Q}^{\prime}|\leq 2|\mathcal{Q}|$ and that the paths of $\mathcal{Q}^{\prime}$ cover all the edges covered by the paths of $\mathcal{Q}$, i.e., such that
$\bigcup_{P\in\mathcal{Q}}E(P)\subseteq\bigcup_{P\in\mathcal{Q}^{\prime}}E(P).$
Proof.
Assume that $\mathcal{Q}$ is not reduced, and let $Q$ be a reducing path in $\mathcal{Q}$. We also let $P$ be a subpath of $Q$ with endvertices $u,v$ such that $u<_{1}v$, and $R$ be a shortest path in $G$ from $u$ to $v$ intersecting $P_{1}$. We let $R_{1},R_{2}$ denote two shortest paths respectively from $a_{1}$ to $u$ and from $v$ to $b_{1}$. We also write $Q=Q_{1}\cdot Q_{2}\cdot Q_{3}$, so that $Q_{2}$ is the $uv$-subpath of $Q$ parallel to $P_{1}$ (see Figure˜2). Observe now that both paths $Q^{\prime}:=Q_{1}\cdot R\cdot Q_{3}$ and $Q^{\prime\prime}:=R_{1}\cdot Q_{2}\cdot R_{2}$ are shortest paths in $G$ that both intersect $P_{1}$ and cover all the edges of $Q$. In particular, it implies that the family of paths $\mathcal{Q}^{\prime}:=(\mathcal{Q}\setminus\{Q\})\cup\{Q^{\prime},Q^{\prime\prime}\}$ covers all the edges covered by the paths of $\mathcal{Q}$, and has strictly less paths than $\mathcal{Q}$ that do not intersect $P_{1}$. We apply iteratively the same operation as long as the obtained family $\mathcal{Q}^{\prime}$ is not reduced. In particular, as at each step, the number of paths of $\mathcal{Q}$ which are disjoint with $P_{1}$ strictly decreases, then after at most $|\mathcal{Q}|$ iterations, we obtain a reduced family with the desired properties. ∎
Figure 2: The path $Q$ represented in green is reducing. $P_{1}$ is represented in blue.
Up to applying Proposition˜3.5 to $\mathcal{P}$, we will assume from now on that $\mathcal{P}$ is a reduced family of at most $2k$ shortest paths of $G$ that edge-cover $G$.
Good and bad vertices.
We now partition the set $\mathcal{P}:=\{P_{i}:1\leq i\leq k\}$ into the subset $\mathcal{P}_{1}$ of the paths in $\mathcal{P}$ that intersect $P_{1}$, and $\mathcal{P}_{2}:=\mathcal{P}\setminus\mathcal{P}_{1}$.
For every $P\in\mathcal{P}_{1}$, we let $C_{P}$ denote the maximal subpath of $P$ with both endvertices $a_{P},b_{P}$ in $V(P_{1})$, so that $a_{P}\leq_{1}b_{P}$ (up to considering $P^{-1}$ instead of $P$, we also assume that $a_{P}$ is before $b_{P}$ on $P$). Note that $C_{P_{1}}=P_{1}$, and that by definition of $\mathcal{P}_{1}$, each $C_{P}$ has at least one vertex, and is parallel to $P_{1}$. We moreover consider for every $P\in\mathcal{P}_{1}$ the two vertices $u_{P},v_{P}$ that respectively maximize $d_{G}(u_{P},a_{P})$ and $d_{G}(b_{P},v_{P})$, such that $u_{P},a_{P},b_{P},v_{P}$ appear in this order on $P$, and such that the paths $P[u_{P},a_{P}]$ and $P[b_{P},v_{P}]$ are parallel to $P_{1}$. We now let $\widetilde{C}_{P}:=P[u_{P},v_{P}]$ and $A_{P},B_{P}$ denote the two subpaths of $P$ such that $P=A_{P}\cdot\widetilde{C}_{P}\cdot B_{P}$ (see Figure˜3). We stress out that in general, the path $\widetilde{C}_{P}$ is not necessarily parallel to $P_{1}$.
Figure 3: The red path represents some path $P\in\mathcal{P}_{1}$. The blue path is $P_{1}$.
We let $\mathcal{A}:=\{A^{-1}_{P}:P\in\mathcal{P}_{1}\}\cup\{B_{P}:P\in\mathcal{P}_{1}\}$ and $\mathcal{B}:=\{\widetilde{C}_{P}:P\in\mathcal{P}_{1}\}$. In particular, $P_{1}\in\mathcal{B}$. Note that by definition, every path in $\mathcal{B}$ is the concatenation of at most $3$ paths that are parallel to $P_{1}$, and recall that $|\mathcal{P}|\leq 2k$, hence we have $|\mathcal{B}|\leq 2k$.
We say that the edges of the paths in $\mathcal{A}\cup\mathcal{P}_{2}$ are bad, and we call a vertex $v\in V(G)$ bad if it is the endvertex of a bad edge. We call every vertex which is not bad good, and we let $V_{b},V_{g}$ denote respectively the set of bad and good vertices of $G$. We moreover set $X_{0}:=\{a_{P}:P\in\mathcal{P}_{1}\}\cup\{b_{P}:P\in\mathcal{P}_{1}\}$ and $X_{1}:=\{u_{P}:P\in\mathcal{P}_{1}\}\cup\{v_{P}:P\in\mathcal{P}_{1}\}$. Note that in particular, $|X_{0}|\leq 4k$ and that the only bad vertices that belong to $P_{1}$ are the vertices $a_{P},b_{P}$ for $P\in\mathcal{P}_{1}$, hence $V(P_{1})\cap V_{b}\subseteq X_{0}$.
We now consider the subgraph $G_{0}:=\bigcup_{P\in\mathcal{B}}P$ of $G$. The following remark immediately follows from the definition of $\mathcal{B}$, the fact that $|\mathcal{P}_{1}|\leq|\mathcal{P}|\leq 2k$ and that every path in $\mathcal{B}$ is edge-covered by at most $3$ paths that are parallel to $P_{1}$.
Remark 3.6.
$G_{0}$ is edge-covered by at most $6k$ paths that are parallel to $P_{1}$.
Lemma 3.7.
We have $V_{g}\subseteq V(G_{0})$. Moreover, $G[V_{g}]$ is a subgraph of $G_{0}$.
Proof.
Let $u$ be a good vertex. As $G$ is covered by $\mathcal{P}$, there exists some path $P\in\mathcal{P}$ such that $u\in V(P)$. In particular, as $u$ is a good vertex, we must have $P\in\mathcal{P}_{1}$, and moreover $u$ must be a vertex of $\widetilde{C}_{P}$. It implies that $u\in V(G_{0})$, and thus that $V_{g}\subseteq V(G_{0})$. The ’Moreover’ part follows from the fact that by definition, no bad edge can be incident to a good vertex. In particular, note that every edge from $E(G)\setminus E(G_{0})$ is bad, hence $G[V_{g}]$ is indeed a subgraph of $G_{0}$. ∎
Our goal from now on will be to prove the following lemma.
Lemma 3.8 (Main).
There exists a set $X\subseteq V(G)$ of size at most $720k^{3}+4k$ that separates every vertex on $P_{1}$ from $V_{b}$ in $G$.
Proof of Theorem˜3.1 using Lemma˜3.8.
We let $X$ be given by Lemma˜3.8. In particular, by Lemma˜3.7, every component of GXG-X that contains a vertex of $P_{1}$ induces a subgraph of $G_{0}$. By Lemma˜2.3, Lemma˜3.2 and Remark˜3.6, $G_{0}$ is $6k$-layered, and thus has pathwidth at most $6k$, allowing us to conclude the proof of Theorem˜3.1. ∎
Reduction to a separation problem in the graph $G_{0}$.
We will now show that, in order to prove Lemma˜3.8, it is enough to separate $V(P_{1})$ from $V_{b}$ in the subgraph $G_{0}$ of $G$. We show that Lemma˜3.8 amounts to prove the following key lemma, that we will prove in Section˜3.2.
Lemma 3.9 (Key lemma).
Let $P$ be a path of $G_{0}$ which is parallel to $P_{1}$, and $A\in\mathcal{A}\cup\mathcal{P}_{2}$ which intersects $P$. Then there exists some set $X_{A,P}\subseteq V(G_{0})$ of vertices of size $30k$ such that $X_{A,P}$ separates every vertex of $V(A)\cap V(P)$ from the vertices of $P_{1}$ in $G_{0}$.
Figure 4: Configuration of Lemma˜3.9.
Proof of Lemma˜3.8 using Lemma˜3.9.
Recall that every path of $\mathcal{B}$ is the concatenation of at most $3$ paths parallel to $P_{1}$. We thus consider a set $\mathcal{B}^{\prime}$ of size at most $3|\mathcal{B}|$ of paths parallel to $P_{1}$ that cover the edges of the paths from $\mathcal{B}$. We let $I$ denote the set of all pairs $(A,P)$ such that $A\in\mathcal{A}\cup\mathcal{P}_{2}$, $P\in\mathcal{B}^{\prime}$, and $V(A)\cap V(P)\neq\emptyset$.
We set
$X:=X_{0}\cup\left(\bigcup_{(A,P)\in I}X_{A,P}\right),$
where the sets $X_{A,P}$ are given by Lemma˜3.9. Note that $|\mathcal{A}\cup\mathcal{P}_{2}|\leq 2|\mathcal{P}_{1}|+|\mathcal{P}_{2}|\leq|\mathcal{P}_{1}|+|\mathcal{P}|\leq 4k$ and $|\mathcal{B}^{\prime}|\leq 3|\mathcal{B}|=3|\mathcal{P}_{1}|\leq 6k$, thus $|I|\leq 24k^{2}$ and $|X|\leq 720k^{3}+4k$.
We now show that $X$ separates $V(P_{1})$ from $V_{b}$ in $G$. Suppose, for contradiction, that there exists a path $R$ in GXG-X connecting a vertex $v$ from $V_{b}$ to a vertex $u$ on $P_{1}$. Among those paths, choose a path $R$ with minimum length. Since $V(P_{1})\cap V_{b}\subseteq X_{0}\subseteq X$ and $v\notin V(P_{1})$, $R$ has length at least $1$. Also, since $R$ is a path of minimum length connecting $V_{b}$ to $P_{1}$, every vertex in RvR-v must be good. Since every vertex in RvR-v is good, there is no bad edge in $R$. Hence, as $\mathcal{P}$ edge-covers $G$, all the edges of $R$ must be covered by the paths from $\mathcal{B}$, implying that $R$ is also a path of $G_{0}$.
This implies in particular that $v\in V(G_{0})\cap V_{b}$. Thus by definition of $G_{0}$ and $V_{b}$, there exist some paths $P\in\mathcal{B}$ and $A\in\mathcal{A}\cup\mathcal{P}_{2}$ such that $v\in V(P)\cap V(A)$. Lemma˜3.9 then implies that $R$ intersects $X_{A,P}\subseteq X$, giving a contradiction. ∎
3.2 Proof of Lemma˜3.9
We fix some path $P$ in $G_{0}$ which is parallel to $P_{1}$, and some path $A\in\mathcal{A}\cup\mathcal{P}_{2}$ which intersects $P$. We will in fact prove that one can find a set $X_{A,P}$ of vertices with the desired size that separates in $G_{0}$ all the vertices of $Q$ from $P_{1}$. We let $Q$ denote the minimal subpath of $P$ containing all vertices in $V(A)\cap V(P)$, and let $b,c$ denote the endvertices of $Q$ such that $b\leq_{1}c$. By Remark˜3.6, $G_{0}$ is coverable by at most $6k$ paths which are all parallel to $P_{1}$. In particular, for every $i\geq 0$, the set $V_{i}:=\{v\in V(G_{0}):d_{G}(v,a_{1})=i\}$ contains at most $6k$ vertices, and separates in $G_{0}$ the sets $\bigcup_{j\leq i}V_{j}$ and $\bigcup_{j\geq i}V_{j}$. We also may assume without loss of generality that $Q$ contains at least $2$ vertices, hence $b<_{1}c$ (otherwise, we conclude by choosing $X_{A,P}:=V(Q)=\{b\}$).
For every $v\in V(G_{0})$, we let $\iota(v)$ denote the unique index $i\geq 0$ such that $v\in V_{i}$. For every path $L$ in $G_{0}$ which is parallel to $P_{1}$, if $u,v$ denote the endvertices of $L$ with $u\leq_{1}v$, we define the projection of $L$ on $P_{1}$ as the set $U_{L}:=\{x\in V(P_{1}):\iota(u)<\iota(x)<\iota(v)\}$. We say that $L$ is $\nearrow$-free if there does not exist in $G_{0}$ a path $R$ parallel to $P_{1}$ that connects a vertex $x\in U_{L}$ to a vertex $y\in V(L)$ with $x\leq_{1}y$ (see Figure˜5). Symmetricaly, we say that $L$ is $\searrow$-free if there does not exist in $G_{0}$ a path $R$ parallel to $P_{1}$ that connects a vertex $x\in U_{L}$ to a vertex $y\in V(L)$ with $x\geq_{1}y$. Our main result in this subsection will be Lemma˜3.12, which states that if $Q$ is $\nearrow$-free or $\searrow$-free, then we can separate $Q$ from $P_{1}$ in $G_{0}$ after removing $30k$ vertices. Before proving this, we will first show in the next two lemmas that $Q$ can be covered by at most two subpaths which are each $\nearrow$-free or $\searrow$-free.
Figure 5: The path $L$ (in red) is not $\nearrow$-free.
Lemma 3.10.
If $A\in\mathcal{A}$, then $Q$ is $\nearrow$-free or $\searrow$-free.
Proof.
We assume first that $A=A^{-1}_{S}$ for some $S\in\mathcal{P}_{1}$, and claim that the case where $A=B_{S}$ is symmetric. We moreover assume that $A[u_{S},b]$ is a subpath of $A[u_{S},c]$, as depicted in Figure˜6, and claim that the case where $A[u_{S},c]$ is a subpath of $A[u_{S},b]$ is symmetric. We show that in this case, $Q$ must be $\nearrow$-free. As $P$, and thus $Q$ are parallel to $P_{1}$, note that every vertex $v\in V(Q)$ satisfies $\iota(v)\in[\iota(b),\iota(c)]$. We let $A^{\prime}$ denote the subpath of $S^{-1}$ which starts at $a_{S}$ and contains $A$ as a subpath.
We assume for sake of contradiction that there exists some path $R$ in $G_{0}$ which is parallel to $P_{1}$ in $G$, and which connects a vertex $x\in U_{Q}$ to a vertex $y\in V(Q)$, with $x\leq_{1}y$. We consider two cases according to whether $a_{S}\leq_{1}x$ or not, described in Figure˜6.
If $a_{S}\leq_{1}x$, then by Lemma˜3.4, the path $P_{1}[a_{S},x]\cdot R\cdot Q[y,c]$ is parallel to $P_{1}$. On the other hand, recall that by definition of $u_{S}$, and as $c\neq u_{S}$ (this is because $b\neq c$ and $A[u_{S},c]$ contains $b$), $A^{\prime}[a_{S},c]$ cannot be parallel to $P_{1}$. In particular, $A^{\prime}[a_{S},c]$ must be strictly longer than $P_{1}[a_{S},x]\cdot R\cdot Q[y,c]$, implying a contradiction as $A^{\prime}$ is a shortest path in $G$.
If $a_{S}>_{1}x$, then as $x\in U_{Q}$, $x>_{1}b$. Then, the path $R\cdot Q[y,c]$ is vertical with respect to $a_{1}$, and thus strictly shorter than $Q$. Note in this case, the path $A^{\prime}[a_{S},b]$ is also strictly longer than $P_{1}[a_{S},x]$. It thus implies that $P_{1}[a_{S},x]\cdot R\cdot Q[y,c]$ is strictly shorter than the path $A^{\prime}[a_{S},c]$, contradicting again that $A^{\prime}$ is a shortest path in $G$. ∎
Figure 6: Top: configuration in the proof of Lemma˜3.10 when $a_{S}\leq_{1}x$. Bottom: configuration in the proof of Lemma˜3.10 when $a_{S}>_{1}x$.
Lemma 3.11.
If $A\in\mathcal{P}_{2}$, then there exists subpaths $Q_{1},Q_{2}$ such that $Q=Q_{1}\cdot Q_{2}$, $Q_{1}$ is $\searrow$-free and $Q_{2}$ is $\nearrow$-free.
Proof.
If $Q$ is $\nearrow$-free, then we immediately conclude by choosing $Q_{1}$ to be the path having $b$ as a single vertex, thus we may assume that there exists some path $R$ parallel to $P_{1}$ connecting some vertex $x\in U_{Q}$ to a vertex $y\in V(Q)$ such that $x\leq_{1}y$. We moreover choose such a path so that $x$ is maximal with respect to $\leq_{1}$. We let $z$ be the unique vertex of $Q$ such that $\iota(z)=\iota(x)$, and set $Q_{1}:=Q[b,z],Q_{2}:=Q[z,c]$. By maximality of $x$, note that $Q_{2}$ must be $\nearrow$-free. We moreover claim that $Q_{1}$ is $\searrow$-free, as otherwise we would be able to find a shortest path $T$ in $G$ from $b$ to $c$ that intersects $P_{1}$ (see Figure˜7). In particular, as $b,c\in V(A)$ the existence of such a path would imply that $A$ is a reducing path belonging to $\mathcal{P}$, contradicting our assumption that $\mathcal{P}$ is reduced. ∎
Figure 7: The configuration at the end of the proof of Lemma˜3.11. The purple dashed path $R^{\prime}$ represents a shortest path path which is parallel to $P_{1}$, and connects vertices $y^{\prime}\leq_{1}x^{\prime}$, with $y^{\prime}\in V(Q_{1})$ and $x^{\prime}\in U_{Q_{1}}$.
The next lemma is the crucial part from the proof of Lemma˜3.9.
Lemma 3.12.
Let $L$ be a path in $G_{0}$ that is parallel to $P_{1}$. If $L$ is $\nearrow$-free or $\searrow$-free, then there exists $X\subseteq V(G_{0})$ of size at most $18k$ that separates in $G_{0}$ the vertices of $L$ from the ones of $P_{1}$.
Proof.
We let $b_{L},c_{L}$ denote the endvertices of $L$, so that $b_{L}\leq_{1}c_{L}$, and let $X^{\prime}_{0}:=V_{\iota(b_{L})}\cup V_{\iota(c_{L})}$. Recall that by Remark˜3.6, $G_{0}$ is edge-coverable by $k^{\prime}$ paths $Q_{1},\ldots,Q_{k^{\prime}}$, all of them parallel to $P_{1}$ for some $k^{\prime}\leq 6k$. In particular, it implies that $X^{\prime}_{0}$ has size at most $12k$, and separates $L$ from every vertex $x$ of $P_{1}$ such that $x\leq_{1}b$ or $x\geq_{1}c$. It remains to separate $L$ from $U_{L}$. We assume without loss of generality that $L$ is $\nearrow$-free, the other case being symmetric, up to exchanging the roles of $a_{1}$ and $b_{1}$. We set $Q_{0}:=L$, and for each $j\in\{0,\ldots,k^{\prime}\}$, we let $H_{j}:=P_{1}\cup Q_{0}\cup Q_{1}\cup\cdots\cup Q_{j}$. Observe that in particular, $H_{0}=P_{1}\cup L$ and $H_{k^{\prime}}=G_{0}$.
Informally, our proof works as follows: the set $X$ separating $P_{1}$ from $L$ that we construct will be the union of $X^{\prime}_{0}$ and of at most one vertex $y_{j}$ for each path $Q_{j}$ ($j\in\{1,\ldots,k^{\prime}\}$). We will choose the vertices $y_{j}$ as follows: we start considering the rightmost (i.e. maximal with respect to $\leq_{1}$) vertex $y_{1}$ which is in the intersection of $L$ and of some path $Q_{j_{1}}$ with $j_{1}\in\{1,\ldots,k^{\prime}\}$. Without loss of generality, we may assume that $j_{1}=1$. Note that by maximality of $y_{1}$, the set $X^{\prime}_{0}\cup\{y_{1}\}$ separates the vertices of $L[y_{1},c_{L}]$ from the ones of $P_{1}$ in $G_{0}$ (and thus also in $H_{1}$). Moreover, as $L$ is $\nearrow$-free, none of the vertices from $L\cup Q_{1}$ which are smaller or equal to $y_{1}$ with respect to $\leq_{1}$ can belong to $P_{1}$, hence $X^{\prime}_{0}\cup\{y_{1}\}$ separates the vertices of $L\cup Q_{1}$ from the ones of $P_{1}$ in $H_{1}$. In particular, observe that if none of the paths $Q_{2},\ldots,Q_{k^{\prime}}$ contains a vertex from $L\cup Q_{1}$ which is smaller or equal to $y_{1}$ with respect to $\leq_{1}$, then it means that $X^{\prime}_{0}\cup\{y_{1}\}$ separates $L$ from $P_{1}$ in $G$. We may thus assume that there exists some vertex $y_{2}\leq_{1}y_{1}$ that belongs to $V(Q_{j_{2}})\cap V(Q_{1}\cup L)$ for some $j_{2}\geq 2$. Again, we choose such a vertex $y_{2}$ maximal with respect to $\leq_{1}$, and assume without loss of generality that $j_{2}=2$. By maximality of $y_{2}$, the set $X^{\prime}_{0}\cup\{y_{1},y_{2}\}$ separates the vertices of $L$ which are greater or equal to $y_{2}$ from the ones of $P_{1}$ in $H_{2}$, and as $L$ is $\nearrow$-free, one can check that none of the vertices from $L\cup Q_{1}\cup Q_{2}$ which are smaller than $y_{2}$ with respect to $\leq_{1}$ can belong to $P_{1}$. We then repeat the construction inductively, by choosing at each step some vertex $y_{j+1}\leq_{1}y_{j}$ maximal with respect to $\leq_{1}$ that belongs to $V(L\cup Q_{1}\cup\cdots\cup Q_{j})\cap V(Q_{i_{j+1}})$ for some $i_{j+1}\geq j+1$ (without loss of generality, we then assume that $i_{j+1}=j+1$), and claim that at each step, the set $X^{\prime}_{0}\cup\{y_{1},\ldots,y_{j+1}\}$ separates the vertices of $L$ from the ones of $P_{1}$ in $H_{j+1}$, and that none of the vertices from $L\cup Q_{1}\cup\cdots\cup Q_{j+1}$ which are smaller than $y_{j+1}$ with respect to $\leq_{1}$ can belong to $P_{1}$. See Figure˜8 for an example.
Figure 8: In the following example, as $Q_{2}$ and $Q_{3}$ both have their rightmost intersection with $W_{1}$ at the same vertex, we will have $y_{2}=y_{3}$, and we can choose $Q_{2}$ and $Q_{3}$ in any order in the proof of Lemma˜3.12.
We set $y_{0}:=c_{L}$ and $W_{0}:=V(L)$. We construct a sequence of vertices $y_{\ell}\leq_{1}\cdots\leq_{1}y_{0}$ and a sequence of sets $W_{0},\ldots,W_{\ell}$ for some $0\leq\ell\leq k^{\prime}$. These sequences are constructed by induction as follows. Let $j\geq 0$ and assume that we already constructed $y_{0},\ldots,y_{j}$ and $W_{0},\ldots,W_{j}$ for some $j<k^{\prime}$, such that, up to permuting the indices of the paths $Q_{1},\ldots,Q_{k^{\prime}}$, for every $j^{\prime}\leq j$, $y_{j^{\prime}}\in V(Q_{j^{\prime}})$. We now let $y_{j+1}$ be any vertex which is maximal with respect to $\leq_{1}$ such that
  • $y_{j+1}\leq_{1}y_{j}$,
  • $y_{j+1}\in W_{j}\cap V\left(\bigcup_{s\geq j+1}Q_{s}\right)$.
If no such vertex exists, then we stop the construction here and set $\ell:=j$. Otherwise, we let $y_{j+1}$ be such a vertex, and, up to permuting the indices of the paths $Q_{j+1},\ldots,Q_{k^{\prime}}$, we assume that $y_{j+1}\in V(Q_{j+1})$. We write $Q_{j+1}=Q_{j+1}^{-}\cdot Q_{j+1}^{+}$, where $Q_{j+1}^{-}$ (resp. $Q_{j+1}^{+}$) denote the subpath of $Q_{j+1}$ containing the vertices which are smaller or equal (resp. greater or equal) to $y_{j+1}$ with respect to $\leq_{1}$. Note that $Q_{j+1}^{-}$ and $Q_{j+1}^{+}$ are internally disjoint and share $y_{j+1}$ as an endvertex. We set $W_{j+1}:=W_{j}\cup V(Q_{j+1}^{-})$. See Figure˜8 for an illustration.
We prove by induction over $j\geq 0$ the following properties.
  1. (i)
    there exists a path $R_{j}$ from $y_{j}$ to $c_{L}$ in $G_{0}$ which is parallel to $P_{1}$;
  2. (ii)
    for every $y\in W_{j}$, if $y>_{1}y_{j}$, then $y\notin V\left(\bigcup_{s\geq j}Q_{s}\right)$;
  3. (iii)
    for every $j^{\prime}\leq j$, $V(Q_{j^{\prime}}^{+})\cap W_{j}\subseteq\{y_{j^{\prime}}\}$;
  4. (iv)
    $W_{j}\cap U_{L}=\emptyset$;
  5. (v)
    $X^{\prime}_{0}\cup\{y_{1},\ldots,y_{j}\}$ separates $W_{j}$ from $U_{L}$ in $H_{j}$.
Note that when $j=0$, items (i) to (v) are immediate (note that (iv) holds in this case because $L$ is $\nearrow$-free, and thus, $P_{1}$ must be internally disjoint from $L$). We now assume that $0\leq j<k^{\prime}$, and that items (i) to (v) hold for any $j^{\prime}\leq j$, and we prove that they are also satisfied for $j+1$.
We start showing (i). First, as $y_{j+1}\in W_{j}$, there exists some $j^{\prime}\leq j$ such that $y_{j+1}\in V(Q_{j^{\prime}})$ and $y_{j+1}\leq_{1}y_{j^{\prime}}$. We consider the path $R_{j^{\prime}}$ given by induction hypothesis. In particular, both $y_{j+1}$ and $y_{j^{\prime}}$ belong to $Q_{j^{\prime}}$, hence by Lemma˜3.4, the path $R_{j+1}:=Q_{j^{\prime}}[y_{j+1},y_{j^{\prime}}]\cdot R_{j^{\prime}}$ is parallel to $P_{1}$, and connects $y_{j+1}$ to $c_{L}$ in $G_{0}$. This proves (i).
We now show (ii). Assume for sake of contradiction that there exists $y\in W_{j+1}$ such that $y>_{1}y_{j+1}$ and $y\in V\left(\bigcup_{s\geq j+1}Q_{s}\right)$. By definition of $W_{j+1}$, note that if $y>_{1}y_{j}$, then $y\in W_{j}$. In particular, by induction hypothesis, if $y>_{1}y_{j}$, then $y\notin V\left(\bigcup_{s\geq j}Q_{s}\right)$, and thus $y\notin V\left(\bigcup_{s\geq j+1}Q_{s}\right)$. We thus have $y_{j+1}<_{1}y\leq_{1}y_{j}$. However, we now obtain a contradiction with the maximality of $y_{j+1}$ with respect to $\leq_{1}$, as $y\in W_{j}\cap V\left(\bigcup_{s\geq j+1}Q_{s}\right)$.
Let us now prove (iii), that is, that for every $j^{\prime}\leq j+1$, $V(Q_{j^{\prime}}^{+})\cap W_{j+1}\subseteq\{y_{j^{\prime}}\}$. First, note that if $j^{\prime}\leq j$, as the vertices of $W_{j+1}\setminus W_{j}$ are exactly the internal vertices of $Q_{j+1}^{-}$, and as $y_{j+1}\leq_{1}y_{j^{\prime}}$, then the only possible vertex of $Q_{j^{\prime}}^{+}$ belonging to $W_{j+1}\setminus W_{j}$ is $y_{j^{\prime}}$. In particular, by induction hypothesis we immediately deduce that $V(Q_{j^{\prime}}^{+})\cap W_{j+1}\subseteq\{y_{j^{\prime}}\}$. It remains to check (iii) when $j^{\prime}=j+1$. Note that it follows from (ii), as every vertex of $Q_{j+1}^{+}$ which is distinct from $y_{j+1}$ is greater than $y_{j+1}$ with respect to $_{1}$.
To show (iv), assume for sake of contradiction that there exists a vertex $x$ in $U_{L}\cap W_{j+1}$. By induction hypothesis, $x\notin W_{j}$, hence $x$ is an internal vertex of $Q_{j+1}^{-}$. In particular, $x<_{1}y_{j+1}$, thus, as $Q_{j+1}$ is parallel to $P_{1}$, the subpath $Q_{j+1}[x,y_{j+1}]$ must be parallel to $P_{1}$. We let $R_{j+1}$ be the path from $y_{j+1}$ to $c_{L}$ given by (i). In particular, as $y_{j+1}\leq_{1}c_{L}$, Lemma˜3.4 implies that the path $Q_{j+1}[x,y_{j+1}]\cdot R_{j+1}$ is also parallel to $P_{1}$, and connects $x\in U_{L}$ to $c_{L}\in V(L)$, with $x\leq_{1}c_{L}$, contradicting that $L$ is $\nearrow$-free.
We now prove (v), i.e., that $X^{\prime}_{0}\cup\{y_{1},\ldots,y_{j+1}\}$ separates $W_{j+1}$ from $U_{L}$ in $H_{j+1}$. Assume for sake of contradiction that this is not the case. In particular, as by (iv), $W_{j+1}\cap U_{L}=\emptyset$, there should exist an edge $yy^{\prime}$ in $H_{j+1}-(X^{\prime}_{0}\cup\{y_{1},\ldots,y_{j+1}\})$, with $y\in W_{j+1}$ and $y^{\prime}\notin W_{j+1}$. By definition of $H_{j+1}$ and $W_{j+1}$, as the edge $yy^{\prime}$ does not belong to $W_{j+1}$, it should be an edge of some path $Q_{j^{\prime}}^{+}$, for some $1\leq j^{\prime}\leq j+1$. In particular, (iii) then implies that $y=y_{j^{\prime}}$, a contradiction as $R$ avoids $\{y_{1},\ldots,y_{j^{\prime}}\}$.
We now conclude the proof of Lemma˜3.12. By definition of $\ell$, the sequence $y_{0},\ldots,y_{\ell}$ cannot be extended further, that is, for every $\ell<s\leq k^{\prime}$, $W_{\ell}$ is disjoint from $\{y\in V(Q_{s}):y\leq_{1}y_{\ell}\}$. We set $X:=X^{\prime}_{0}\cup\{y_{1},\ldots,y_{\ell}\}$. By (v), $X$ separates $W_{\ell}$, and thus $L$ from $U_{L}$ in $H_{j+1}$. It remains to show that $X$ also separates $W_{\ell}$ from $U_{L}$ in $G_{0}$. Assume for sake of contradiction that it is not the case. In particular, as $Q_{1},\ldots,Q_{k^{\prime}}$ edge-cover $G_{0}$, there would exist an edge $yy^{\prime}\in E(Q_{s})$ for some $s>\ell$, and some vertices $y\in W_{\ell}$ and $y^{\prime}\in V(G_{0})\setminus W_{\ell}$, such that $b_{L}\leq_{1}y\leq_{1}c_{L}$ and $b_{L}\leq_{1}y^{\prime}\leq_{1}c_{L}$. By our above remark, we have $y>_{1}y_{\ell}$. Note that we then immediately obtain a contradiction with (ii). This implies that $X$ indeed separates $L$ from $P_{1}$ in $G_{0}$. Finally, observe that
$|X|\leq|X^{\prime}_{0}|+k^{\prime}\leq 3k^{\prime}\leq 18k$
Proof of Lemma˜3.9.
By Lemmas˜3.10 and 3.11, there exist two subpaths $Q_{1},Q_{2}$ of $Q$ such that $Q=Q_{1}\cdot Q_{2}$, and such that for each $i\in\{1,2\}$, $Q_{i}$ is either $\nearrow$-free or $\searrow$-free. In particular, by Lemma˜3.12, there exist two sets $X_{1},X_{2}$ of size at most $18k$ such that for each $i\in\{1,2\}$, $X_{i}$ separates in $G_{0}$ the set $V(Q_{i})$ from $V(P_{1})$. In particular, the set $X:=X_{1}\cup X_{2}$ then separates $V(Q)$ from $V(P_{1})$. Moreover, we claim that when going through the proof of Lemma˜3.12, one gets that $X_{1}$ and $X_{2}$ both contain a common layer $V_{i}$ (for $i:=\iota(c_{Q_{1}})=\iota(b_{Q_{2}})$), allowing us to slightly reduce our upper bound on $|X_{1}\cup X_{2}|$ to obtain $|X_{1}\cup X_{2}|\leq 30k$ (instead of $36k$). ∎
4 Exact Bounds for $\boldsymbol{k\leq 3}$
In this section, we prove some optimal upper-bounds on the pathwidth of graphs that are edge-coverable by $k$ shortest paths, when $k\in\{2,3\}$.
See 1.3
4.1 Two Paths
As a warm-up, we start describing the structure of graphs edge-coverable by $2$ shortest paths, implying an immediate proof of Theorem˜1.3 when $k=2$.
Definition 4.1.
A skewer is a graph $G$ for which there exists a sequence of pairwise distinct vertices $x_{1},\ldots,x_{\ell}$ ($\ell\geq 1$) such that:
  • for each $i\in\{1,\ldots,\ell-1\}$, the vertex $x_{i}$ is connected to the vertex $x_{i+1}$ by two internally disjoint paths $Q_{i},R_{i}$ having the same length;
  • for each $i\in\{1,\ldots,\ell-1\}$, the paths $Q_{i},R_{i}$ are separated from $G-(Q_{i}\cup R_{i})$ by the set $\{x_{i},x_{i+1}\}$;
  • for $i\in\{0,\ell\}$, there exist two internally disjoint paths $Q_{i},R_{i}$ (which do not necessarily have the same length), that are separated from $G-(Q_{i}\cup R_{i})$ by the set $\{x_{i}\}$.
See Figure˜9 for an illustration of a skewer. For each $i\in\{0,\ldots,\ell\}$, the graph $C_{i}:=Q_{i}\cup R_{i}$ is called the piece of $x_{i}$. Note that we allow the paths $Q_{i}$ and $R_{i}$ to have length $1$, in which case the graph $C_{i}$ is just the edge $x_{i}x_{i+1}$ (recall that the graphs we consider are simple), while if $Q_{i}$ and $R_{i}$ have length at least $2$, then $C_{i}$ is an even cycle. Note that every skewer $G$ has a $2$-layering, and thus pathwidth at most $2$, and that moreover, for every $i\in\{0,\ldots,\ell\}$, and every pair of vertices $a,b$ with $a\in V(Q_{i})$ and $b\in V(R_{i})$, there exists a path-decomposition of $G$ of width $2$ in which there exists a bag equals to $\{a,b\}$.
Figure 9: A skewer. The blue paths represent the paths $R_{i}$ and the red paths represent the paths $Q_{i}$.
Lemma 4.2.
Let $P,Q$ be two shortest paths in a graph $G$. Then there cannot exist three vertices $a,b,c\in V(P)\cap V(Q)$ such that $b$ is an internal vertex of $P[a,c]$ and such that $c$ is an internal vertex of $Q[a,b]$.
Proof.
Assume for the sake of contradiction that there exist three such vertices $a,b,c\in V(P)\cap V(Q)$. Since $b$ is an internal vertex of $P[a,c]$, $d(a,b)<d(a,c)$. But since $c$ is an internal vertex of $Q[a,b]$, $d(a,c)<d(a,b)$, a contradiction. ∎
Proposition 4.3.
Every connected graph $G$ which is edge-coverable by two isometric paths is a skewer. In particular, $\mathrm{pw}(G)\leq 2$.
Proof.
We let $P_{1},P_{2}$ be two isometric paths that edge-cover $G$, and let $x_{1},\ldots,x_{\ell}$ denote the vertices from $V(P_{1})\cap V(P_{2})$, such that for each $i<j$, $x_{i}$ is before $x_{j}$ on $P_{1}$. By Lemma˜4.2, for each $i<j$, $x_{i}$ is also before $x_{j}$ on $P_{2}$. It thus implies that each $x_{i}$ separates in $G$ all vertices appearing before it in $P_{1}$ and $P_{2}$ from all vertices appearing after it in $P_{1}$ and $P_{2}$. In particular, it easily follows that $G$ is a skewer. ∎
4.2 Three Paths
This subsection consists in a proof of Theorem˜1.3 when $k=3$, which we restate here for convenience.
Theorem 4.4.
Let $G$ be a graph which is edge-coverable by $3$ isometric paths. Then $\mathrm{pw}\left(G\right)\leq 3$.
In the remainder of the subsection, we let $G$ denote a graph which is edge-coverable by three shortest paths $P_{1},P_{2},P_{3}$. We also assume that for each $i\in\{1,2,3\}$, $P_{i}$ is a $a_{i}b_{i}$-path for some vertices $a_{i},b_{i}\in V(G)$. Moreover, note that if the $P_{i}$’s do not all belong to the same connected component of $G$, then every connected component of $G$ is coverable by at most $2$ shortest paths, hence by Proposition˜4.3, $\mathrm{pw}(G)\leq 2$. We thus moreover assume that $G$ is connected.
[…]
Lemma 4.5.
If the intersection of two of the paths $P_{1},P_{2},P_{3}$ is included in the third one, then $\mathrm{pw}(G)\leq 3$.
Proof.
Assume without loss of generality that $V(P_{1})\cap V(P_{2})\subseteq V(P_{3})$, and that $G$ is connected. Note that if one of the two paths $P_{1},P_{2}$ is disjoint from $P_{3}$, then $G$ is not connected and we get a contradiction, hence both paths intersect $P_{3}$. For each $i\in\{1,2\}$, we let $u_{i},v_{i}$ denote the two vertices from $V(P_{i})\cap V(P_{3})$ maximizing $d(u_{i},v_{i})$, and such that $d(a_{3},u_{i})\leq d(a_{3},v_{i})$, and we let $B_{i}:=P_{i}[u_{i},b_{i}]$ (see Figure˜10). Then the graph $H:=P_{3}\cup B_{1}\cup B_{2}$ admits a $3$-layering $(V_{0},\ldots,V_{r})$, where $r:=\mathrm{ecc}_{H}(a_{3})$, and for each $i\in\{0,\ldots,r\}$, $V_{i}:=\{u\in V(H):d(a_{3},u)=i\}$; this follows from the observation that the three paths $P_{3},B_{1},B_{2}$ are vertical with respect to $a_{3}$ (note that the hypothesis that $V(P_{3})\subseteq V(P_{1})\cap V(P_{2})$ is used here). In particular, note that this layering satisfies the property that each path $P_{j}$ intersects at most once each layer $V_{i}$. Observe now that $G$ is obtained from $H$ by attaching the two pendant paths $A_{1}:=P_{1}[a_{1},u_{1}]$ and $A_{2}:=P_{2}[a_{2},u_{2}]$ to the vertices $u_{1}$ and $u_{2}$. We claim that it is then not hard to construct a $3$-layering of $G$ obtained from the one of $H$ by adding these two pendant paths to the layering $(V_{i})_{0\leq i\leq r}$. ∎
Figure 10: Proof of Lemma˜4.5.
Lemma 4.6.
If one of the three paths $P_{1},P_{2},P_{3}$ intersects the other two at most twice in total, then $\mathrm{pw}(G)\leq 3$.
Proof.
Assume without loss of generality that $|V(P_{3})\cap(V(P_{1})\cup V(P_{2}))|\leq 2$. Note that if $P_{3}$ does not intersect $V(P_{1})\setminus V(P_{2})$ (or symmetrically, $V(P_{2})\setminus V(P_{1})$), then we have $V(P_{3})\cap V(P_{1})\subseteq V(P_{2})$ (and symmetrically, $V(P_{3})\cap V(P_{2})\subseteq V(P_{1})$), in which case Lemma˜4.5 implies that $\mathrm{pw}(G)\leq 3$. We may thus assume that $P_{3}$ intersects $V(P_{1})\cup V(P_{2})$ in exactly two vertices $a\in V(P_{1})\setminus V(P_{2})$ and $b\in V(P_{2})\setminus V(P_{1})$.
We now set $H:=P_{1}\cup P_{2}$. By Proposition˜4.3, $H$ is a skewer, and we let $x_{1},\ldots,x_{\ell}$, $R_{0},\ldots,R_{\ell}$, $Q_{0},\ldots,Q_{\ell}$, denote its corresponding components with the notations of Definition˜4.1. Assume first that there exists some $i\in\{0,\ell\}$, such that $a\in R_{i}$ and $b\in Q_{i}$. Then $H$ has a path-decomposition of width $2$ that contains a bag equal to $\{a,b\}$. In particular, as $\{a,b\}$ separates $V(P_{3})$ from $V(H)$ in $G$, we can insert after the bag $\{a,b\}$ of this path-decomposition some path-decomposition of width $1$ of $P_{3}$ to which we add in each bag the set $\{a,b\}$. This gives a path-decomposition of width at most $3$ of $G$.
Assume now that we are not in the previous situation. In particular, as we assumed that $a\in V(P_{1})\setminus V(P_{2})$ and $b\in V(P_{2})\setminus V(P_{1})$, up to symmetry, there exist $i\neq j$ such that $a\in R_{i}$ and $b\in Q_{j}$. We now let $P^{\prime}_{1}:=R_{0}\cdots R_{j-1}\cdot Q_{j}\cdot R_{j+1}\cdots R_{\ell}$ and $P^{\prime}_{2}:=Q_{0}\cdots Q_{j-1}\cdot R_{j}\cdot Q_{j+1}\cdots Q_{\ell}$. Observe that $H$ is edge-covered by $P^{\prime}_{1},P^{\prime}_{2}$, and that $P^{\prime}_{1},P^{\prime}_{2}$ are shortest paths in $G$, such that $a,b\in V(P^{\prime}_{1})$. In particular, we can again conclude as above, using Lemma˜4.5 with respect to the paths $P^{\prime}_{1},P^{\prime}_{2},P_{3}$ that edge-cover $G$. ∎
Lemma 4.7.
If there are at least two vertices that simultaneously belong to the three paths $P_{1},P_{2},P_{3}$, then $\mathrm{pw}(G)\leq 3$
Proof.
Let $u$ and $v$ be two distinct vertices of $V(P_{1})\cap V(P_{2})\cap V(P_{3})$. Without loss of generality, we assume that $a_{i},u,v,b_{i}$ appear in this order on $P_{i}$ for $i=1,2,3$. For $i=1,2,3$, we denote $A_{i}$ the $a_{i}u$-subpath of $P_{i}$, and $B_{i}$ the $ub_{i}$-subpath of $P_{i}$. Then, the graph $G$ admits a $3$-layering $(U_{\ell},\dots,U_{1},V_{0},\dots,V_{r})$ such that $U_{i}:=\{x\in V(A_{1})\cup V(A_{2})\cup V(A_{3}):d(u,x)=i\}$ and $V_{i}:=\{x\in V(B_{1})\cup V(B_{2})\cup V(B_{3}):d_{H}(u,x)=i\}$; this follows from the fact that $\{u\}$ separates the vertices on the $A_{i}$’s from the vertices on the $B_{i}$’s in $G$, and that the $A_{i}$’s and $B_{i}$’s are all vertical w.r.t. $u$. ∎
We say that a path $P$ bounces between two paths $Q_{1},Q_{2}$ if there exists $i\in\{1,2\}$ such that $P$ has a subpath $P^{\prime}$ with both endvertices on $Q_{i}$, and such that one of the internal vertices of $P^{\prime}$ intersects $Q_{3-i}$.
By Lemma˜4.6, we can assume from now on that $P_{1}$ intersects the other paths at least twice, and by Lemma˜4.7, that at most one of the vertices in these intersections lies simultaneously on the three paths $P_{1},P_{2},P_{3}$. In particular, up to exchanging $a_{1}$ and $b_{1}$, and $P_{2}$ and $P_{3}$, we may assume that if $u_{0}$ denotes the vertex on $P_{1}$ which is the closest to $a_{1}$, and belongs to one of the two paths $P_{2},P_{3}$, then $u_{0}\in V(P_{1})\cap V(P_{2})\setminus V(P_{3})$. We set $A_{i}:=P_{i}[a_{i},u_{0}]$ and $B_{i}:=P_{i}[u_{0},b_{i}]$ for $i=1,2$.
Remark 4.8.
Observe that our choice of $u_{0}$ implies that $V(A_{1})\cap(V(P_{2})\cup V(P_{3}))=\{u_{0}\}$.
By Lemma˜4.5 and Lemma˜4.6, we can assume that $P_{3}$ intersects $B_{1}-u_{0}$, and at least one of the paths between $A_{2}-u_{0}$ and $B_{2}-u_{0}$, say $B_{2}-u_{0}$ by symmetry. In particular, at least one of the following situations occurs (up to symmetrically exchange the roles of $A_{2}$ and $B_{2}$)
  1. (1)
    $P_{3}$ bounces between two of the paths $A_{2},B_{2}$ and $B_{1}$,
  2. (2)
    $P_{3}$ does not intersect $A_{2}$,
  3. (3)
    $P_{3}$ intersects the three paths $A_{2},B_{2},B_{1}$ in this order,
  4. (4)
    $P_{3}$ intersects the three paths $A_{2},B_{1},B_{2}$ in this order.
To conclude the proof of Theorem˜1.3, we will prove that $\mathrm{pw}(G)\leq 3$ in each of these cases in the following separate lemmas. This is the most technical part of the proof, especially for cases (1) and (2). The next four lemmas respectively deal with each of the four cases. We start with (1).
Lemma 4.9.
If $P_{3}$ bounces between two of the paths in $\{A_{2},B_{1},B_{2}\}$, then $\mathrm{pw}(G)\leq 3$.
Proof.
We let $u_{0}\in V(P_{1})\cap V(P_{2})\setminus V(P_{3})$ be such that $P_{3}$ bounces between two paths in $\{A_{2},B_{1},B_{2}\}$. Note that by Claim 4.2, $P_{3}$ cannot bounce between $A_{2}$ and $B_{2}$. We assume that $P_{3}$ bounces between $B_{1}$ and $B_{2}$, and claim that the case where $P_{3}$ bounces between $B_{1}$ and $A_{2}$ is symmetric. Let $B_{3}$ be the longest subpath of $P_{3}$ with both extremities $u_{1}$ and $u_{3}$ in $V(B_{1})\cup V(B_{2})$. Since $P_{3}$ bounces between $B_{1}$ and $B_{2}$, $B_{3}$ has length at least $2$. We moreover assume without loss of generality that $0<d(u_{0},u_{1})\leq d(u_{0},u_{3})$ (recall that $u_{0}\notin V(P_{3})$).
We claim that $B_{3}$ is vertical with respect to $u_{0}$. This is immediate if $B_{3}$ has both endvertices on $B_{1}$, or symmetrically on $B_{2}$, as both $B_{1},B_{2}$ are shortest paths in $G$ starting from $u_{0}$. Assume now without loss of generality that $u_{1}\in V(B_{1})$, while $u_{3}\in V(B_{2})$. Then one can easily see that, as $B_{3}$ bounces between $B_{1}$ and $B_{2}$, there exist two distinct vertices $u_{2}\in V(B_{2}),u^{\prime}_{2}\in V(B_{1})$ such that $u_{1},u_{2},u^{\prime}_{2},u_{3}$ are pairwise distinct and appear in this order on $B_{3}$. In particular, in the latter case, $B_{3}[u_{1},u^{\prime}_{2}]$ and $B_{3}[u_{2},u_{3}]$ both bounce between $B_{1}$ and $B_{2}$, and have both endvertices respectively on $B_{1}$, and on $B_{2}$, so they are both vertical with respect to $u_{0}$. As they edge-cover $B_{3}$, and share the non-trivial proper subpath $B_{3}[u_{2},u^{\prime}_{2}]$, it follows that $B_{3}$ is also vertical with respect to $u_{0}$.
We now prove that $B_{3}$ does not intersect $A_{2}$. Assume for sake of contradiction that there exists a vertex $w$ in $V(B_{3})\cap V(A_{2})$. As $u_{0}\notin V(P_{3})$, $w\neq u_{0}$. As $B_{3}$ bounces between $B_{1}$ and $B_{2}$, note that $V(B_{3})\cap V(B_{2})$ contains a vertex $v$ distinct from $u_{0}$. By the previous paragraph, observe that the path $Q:=B_{3}[v,w]$ is vertical with respect to $u_{0}$, hence $d(v,w)<\max(d(u_{0},v),d(u_{0},w))$. In particular, we then obtain a contradiction as $P_{2}[v,w]$ has length $d(u_{0},v)+d(u_{0},w)$, and should be a shortest $vw$-path in $G$.
We now consider the subpaths $A_{3},C_{3}$ of $P_{3}$ such that $P_{3}:=A_{3}\cdot B_{3}\cdot C_{3}$. By construction of $B_{3}$, both $A_{3}-u_{1}$ and $C_{3}-u_{3}$ are disjoint from $V(B_{1})\cup V(B_{2})$.
We prove that $C_{3}$ does not intersect $A_{2}$. Assume for sake of contradiction that $C_{3}$ intersects $A_{2}$, on a vertex $w$ (see Figure˜11). First, note that $B_{3}$ should then intersect $B_{2}$ in at most one vertex, as otherwise, the paths $P_{2},P_{3}$ would contradict Lemma˜4.2. We then assume that $B_{3}$ intersects $B_{2}$ in exactly one vertex $u_{2}$. Since $B_{3}$ bounces between $B_{1}$ and $B_{2}$, in particular $u_{2}$ is an internal vertex of $B_{3}$, and we must have $u_{1},u_{3}\in V(B_{1})\setminus V(B_{2})$. We now set $a:=d(u_{1},u_{2})$, $b:=d(u_{2},u_{3}),c:=d(u_{3},w),d:=d(u_{0},w),e:=d(u_{0},u_{2})$ and $f:=d(u_{0},u_{1})$, so that the situation is as depicted in Figure˜11. Since $B_{3}$ is vertical with respect to $u_{0}$, we have $e=a+f$. Moreover, since both paths $P_{2}[w,u_{2}]$ and $P_{3}[u_{2},w]$ are shortest paths, we have $d(u_{2},w)=b+c=e+d$. Moreover, as $P_{3}[u_{1},w]$ is a shortest path, its length is shorter than the one of $P_{2}[u_{0},w]\cdot P_{1}[u_{0},u_{1}]$, giving $a+b+c\leq f+d$. Combining all previous inequalities then gives $a+e\leq f$, and thus $a=0$, contradicting that $u_{2}$ is an internal vertex of $B_{3}$. We then obtain, as desired, that $C_{3}$ does not intersect $A_{2}$, and thus that $C_{3}$ only intersects $P_{1},P_{2}$ in $u_{3}$.
Figure 11: Configuration from the proof of Lemma˜4.9. The dotted lines represent the layers in the BFS from $u_{0}$.
We now prove that $V(B_{1})\cap V(A_{2})=\{u_{0}\}$. Assume for sake of contradiction that there exists a vertex $v\in V(B_{1})\cap V(A_{2})$ distinct from $u_{0}$. As $B_{3}$ bounces between $B_{1}$ and $B_{2}$ and as $u_{0}\notin V(P_{3})$, there exist two distinct vertices $x,y$ of $B_{3}$ such that $0<d(x,u_{0})<d(y,u_{0})$, and such that $x\in V(B_{1})$ and $y\in V(B_{2})$. We distinguish two cases (depicted in Figure˜12), according whether $d(u_{0},x)\leq d(u_{0},v)$ or not. Note that in both cases, as $A_{2},B_{1},B_{2}$ and $B_{3}$ are vertical with respect to $u_{0}$, and as $x,v\neq u_{0}$, we obtain a contradiction, as the path $B_{1}[v,x]\cdot B_{3}[x,y]$ is always strictly shorter than the path $P_{2}[v,y]$, which is supposed to be a shortest path in $G$.
Figure 12: The two different cases considered in the proof of Lemma˜4.9, when we assume that $A_{2}$ and $B_{2}$ intersect each other. On the left, the situation where $d(u_{0},x)\leq d(u_{0},v)$, and on the right, the situation where $d(u_{0},x)\geq d(u_{0},v)$.
All previous paragraphs together with Remark˜4.8 thus imply that we can write $G=G_{1}\cup G_{2}$, where $G_{1}:=A_{1}\cup A_{2}\cup A_{3}$ and $G_{2}:=B_{1}\cup B_{2}\cup B_{3}\cup C_{3}$, and $\{u_{0},u_{1}\}$ separates $G_{1}$ from $G_{2}$. We now show that $G_{1}$ (resp. $G_{2}$) admits a path-decomposition of width at most $3$ whose rightmost (resp. leftmost) bag contains $\{u_{0},u_{1}\}$. This will immediately imply that $\mathrm{pw}(G)\leq 3$.
We first construct such a path-decomposition $\mathcal{P}_{2}$ for $G_{2}$. Observe that $B_{1},B_{2}$ are both vertical with respect to $u_{0}$. Moreover, recall that $C_{3}-u_{3}$ is disjoint from $V(B_{1})\cup V(B_{2})$, hence, as $B_{3}$ is vertical with respect to $u_{0}$, it implies that $B_{3}\cdot C_{3}$ is also vertical with respect to $u_{0}$. It thus implies that the partition $(V_{i})_{i\geq 0}$ of $V(G_{2})$ defined by setting for each $i\in\{0,\ldots,r\}$ (where $r:=\mathrm{ecc}_{G_{2}}(u_{0})$), $V_{i}:=\{x\in V(G_{2}):d(u_{0},x)=i\}$ is a $3$-layering of $G_{2}$. Moreover, note that if we set $s:=d(u_{0},u_{1})$, then the two subsequences $\mathcal{S}_{1}:=(V_{0},\ldots,V_{s})$ and $\mathcal{S}_{2}:=(V_{s},\ldots,V_{r})$ of $(V_{i})_{0\leq i\leq r}$ are respectively $2$- and $3$-layerings of the subgraphs $G^{\prime}_{2},G^{\prime\prime}_{2}$ of $G_{2}$ induced respectively by $V_{0}\cup\cdots\cup V_{s}$ and by $V_{s}\cup\cdots\cup V_{r}$ (see Figure˜13). In particular, two applications of Lemma˜2.3 give path-decompositions $\mathcal{P}^{\prime}_{2},\mathcal{P}^{\prime\prime}_{2}$ respectively of $G^{\prime}_{2}$ and $G^{\prime\prime}_{2}$ such that:
  • $\mathcal{P}^{\prime}_{2}$ has width at most $2$, and its extremities are $V_{0}=\{u_{0}\}$ and $V_{s}$,
  • $\mathcal{P}^{\prime\prime}_{2}$ has width at most $3$, and its leftmost bag is $V_{s}$ (and thus contains $u_{1}$).
To obtain a path-decomposition $\mathcal{P}_{2}$ of $G_{2}$ with the desired properties, we can now simply add the vertex $u_{1}$ in every bag of $\mathcal{P}^{\prime}_{2}$, and glue the obtained path-decomposition with $\mathcal{P}^{\prime\prime}_{2}$.
Figure 13: The 33-layering of $G_{2}$ described at the end of the proof of Lemma˜4.9.
It remains to construct a path-decomposition $\mathcal{P}_{1}$ of width $3$ of $G_{1}$ whose rightmost bag contains $\{u_{0},u_{1}\}$. By Remark˜4.8, $\{u_{0}\}$ separates the vertices of $A_{1}$ from the other vertices of $G_{1}$. Moreover, by Proposition˜4.3, $A_{2}\cup A_{3}$ is a skewer. In particular, it admits a path-decomposition $\mathcal{P}^{\prime}_{1}$ of width at most $2$ whose rightmost bag is $\{u_{1}\}$. As $G_{1}$ is obtained from $A_{2}\cup A_{3}$ after attaching some pendant path $A_{1}$ on $u_{0}$, the existence of $\mathcal{P}_{1}$ easily follows. ∎
The next lemma deals with (2).
Lemma 4.10.
If $P_{3}$ and $A_{2}$ are disjoint, then $\mathrm{pw}(G)\leq 3$.
Proof.
Suppose that $P_{3}$ and $A_{2}$ are disjoint. Since by Remark˜4.8, $P_{3}$ does not intersect $A_{1}$, by Lemma˜4.5 we may assume that $P_{3}$ intersects both $B_{1}$ and $B_{2}$. We let $v_{1}$ and $w_{1}$ (resp. $v_{2}$ and $w_{2}$) be the vertices of $V(B_{1})\cap V(P_{3})$ (resp. $V(B_{2})\cap V(P_{3})$) such that $v_{1}$ (resp. $v_{2}$) is the closest to $u_{0}$ and $w_{1}$ (resp. $w_{2}$) the farthest. Note that it is possible that $v_{1}=w_{1}$, or $v_{2}=w_{2}$. By Lemma˜4.9, we may assume that $P_{3}$ does not bounce between $B_{1}$ and $B_{2}$, hence we may assume (up to symmetry) that $P_{3}$ intersects $B_{1}$, and then $B_{2}$. By Lemma˜4.7, we may moreover assume that there is at most one vertex in $V(P_{1})\cap V(P_{2})\cap V(P_{3})$. We write $P_{3}=A_{3}\cdot B_{3}\cdot C_{3}$, such that $B_{3}$ is the subpath of $P_{3}$ from the last vertex of $V(P_{1})\cap V(P_{3})$ to the first vertex of $V(P_{2})\cap V(P_{3})$ (note that $B_{3}$ might contain only one vertex if $V(P_{1})\cap V(P_{2})\cap V(P_{3})$ is not empty). By Lemma˜4.2, $B_{3}$ has one endpoint in $\{v_{1},w_{1}\}$, and the other in $\{v_{2},w_{2}\}$. Then, let $C_{1}:=P_{1}[a_{1},v_{1}],D_{1}:=P_{1}[v_{1},w_{1}],E_{1}:=P_{1}[w_{1},b_{1}]$ and let $C_{2}:=P_{2}[a_{2},v_{2}],D_{2}:=P_{2}[v_{2},w_{2}],E_{2}:=P_{2}[w_{2},b_{2}]$ (see Figure˜14).
Figure 14: Configuration in the proof of Lemma˜4.10. In this special case, the endvertices of $B_{3}$ are $v_{2},w_{1}$.
In what follows, we let $c_{1}$ be the first vertex of $P_{3}$ that belongs to $V(P_{1})\cup V(P_{2})$, and $c_{2}$ be the last vertex of $P_{3}$ that belongs to $V(P_{1})\cup V(P_{2})$. As we assumed that $P_{3}$ intersects $B_{1}$ and then $B_{2}$, we have $c_{1}\in V(P_{1})$ and $c_{2}\in V(P_{2})$. Moreover, note that if $c_{1}$ is in $V(P_{2})$, then we have $V(P_{3})\cap V(P_{1})\subseteq V(P_{2})$, and we conclude that $\mathrm{pw}(G)\leq 3$ by Lemma˜4.5. We thus may assume that $c_{1}\in V(P_{1})\cup V(P_{3})\setminus V(P_{2})$, and by symmetry that $c_{2}\in V(P_{2})\cup V(P_{3})\setminus V(P_{1})$.
Claim 4.11.
If one of the two inclusions $V(D_{1})\cap V(B_{2})\subseteq\{v_{1},w_{1}\}$ or $V(D_{2})\cap V(B_{1})\subseteq\{v_{2},w_{2}\}$ does not hold, then $\mathrm{pw}(G)\leq 3$.
Proof of the claim.
We assume that the first inclusion $V(D_{1})\cap V(B_{2})\subseteq\{v_{1},w_{1}\}$ does not hold, and claim that the proof if the second inclusion does not hold is symmetric. Then there exists $v\in V(D_{1})\cap V(B_{2})$ such that $d(u_{0},v_{1})<d(u_{0},v)<d(u_{0},w_{1})$.
We now claim that we are in the same configuration as in Case (1), with $c_{2}$ playing the role of $u_{0}$, where $P_{1}$ bounces between the paths $P_{2}[c_{2},a_{2}]$ and $P_{3}[c_{2},a_{3}]$, as it either intersects $v_{1},v,w_{1}$ in this order, or $w_{1},v,v_{1}$ in this order (see Figure˜15, left). We thus conclude that $\mathrm{pw}(G)\leq 3$ thanks to Lemma˜4.9. ∎
Claim 4.12.
If $V(E_{1})\cap V(C_{2})\neq\emptyset$, or $V(E_{2})\cap V(C_{1})\neq\emptyset$, then $\mathrm{pw}(G)\leq 3$.
Proof of the claim.
Assume that there exists a vertex $v$ in $V(E_{1})\cap V(C_{2})$. We claim that the case where $V(E_{2})\cap V(C_{1})\neq\emptyset$ is symmetric. The situation is then as depicted in the right of Figure˜15. We now set $P^{\prime}_{1}:=A_{1}\cdot P_{2}[u_{0},v]\cdot P_{1}[v,b_{1}]$ and $P^{\prime}_{2}:=A_{2}\cdot P_{1}[u_{0},v]\cdot P_{2}[v,b_{2}]$. Observe that $P^{\prime}_{1},P^{\prime}_{2}$ are also shortest paths in $G$, and that $P^{\prime}_{1},P^{\prime}_{2},P_{3}$ cover the edges of $G$. In particular, note that $V(P_{3})\cap V(P^{\prime}_{1})\subseteq V(P^{\prime}_{2})$, thus Lemma˜4.5 applies and gives $\mathrm{pw}(G)\leq 3$. ∎
Figure 15: Left: configuration of the proof of ˜4.11. In this figure, we depicted the special case where $c_{1}=v_{1}$, however note that there is also a symmetric case where $c_{1}=w_{1}$, in which case $c_{2}$ is before $v$ on $P_{2}$. Right: configuration of the proof of ˜4.12.
Now, observe that Claims 4.11 and 4.12, together with the definition of $v_{1},v_{2}$ imply that we may assume that the set $\{v_{1},v_{2}\}$ separates in $G$ the subgraph $G_{1}:=A_{1}\cdot C_{1}\cup A_{2}\cdot C_{2}$ from the subgraph $G_{2}:=D_{1}\cup E_{1}\cup D_{2}\cup E_{2}\cup P_{3}$ (see Figure˜14). By Proposition˜4.3, $G_{1}$ is a skewer, and admits a path-decomposition $\mathcal{P}^{\prime}_{1}$ of width at most $2$ such that $v_{1}$ is in the rightmost bag of $\mathcal{P}^{\prime}_{1}$. By adding $v_{2}$ to every bag of $\mathcal{P}_{1}^{\prime}$, we obtain a path-decomposition $\mathcal{P}_{1}$ of $G_{1}$ of width at most $3$ such that $v_{1},v_{2}$ is the rightmost bag of $\mathcal{P}_{1}$. It thus remains to show that there exists a path-decomposition of width $3$ of $G_{2}$ whose leftmost bag is $\{v_{1},v_{2}\}$.
We now have four distinct cases, two of them being symmetric, depicted on Figure˜16 according to the order on which the vertices $v_{1},w_{1},v_{2},w_{2}$ appear on $P_{3}$.
Figure 16: Possible ways for the path $P_{3}$ (in dark green) to intersect $P_{1}$ (in blue) and $P_{2}$ (in red) in $G_{2}$. Left: the case where $v_{1},w_{1},w_{2},v_{2}$ appear in this order on $P_{3}$. Middle: the case where $v_{1},w_{1},v_{2},w_{2}$ appear in this order on $P_{3}$. The case where $w_{1},v_{1},w_{2},v_{2}$ appear in this order on $P_{3}$ is symmetric. Right: the case where $w_{1},v_{1},v_{2},w_{2}$ appear in this order on $P_{3}$.
Note that by Claims 4.11 and 4.12, the only possible common vertices between $P_{1}$ and $P_{2}$ in $G_{2}$ are in $V(D_{1})\cap V(D_{2})$, and possibly, the vertices $v_{1},v_{2}$ in the special case when $v_{1}=v_{2}$. In particular, by definition of $w_{1},w_{2}$, the set $\{w_{1},w_{2}\}$ separates in $G_{2}$ the vertices of the graph $G_{3}:=D_{1}\cup D_{2}$ from the remaining vertices. By Proposition˜4.3, and as $u_{0}\in V(P_{1})\cap V(P_{2})$, $G_{3}$ admits a path-decomposition of width at most $2$ whose leftmost bag is $\{w_{1},w_{2}\}$. To conclude our proof, it thus only remains to show in each case that there exists a path-decomposition of $G_{4}:=P_{3}\cup E_{1}\cup E_{2}$ of width at most $3$, whose leftmost bag is $\{v_{1},v_{2}\}$, and whose rightmost bag is $\{w_{1},w_{2}\}$. We claim that it is then not hard to construct separately in each case such a decomposition. ∎
The next two lemmas respectively tackle cases (3) and (4).
Lemma 4.13.
If $P_{3}$ intersects the subpaths $A_{2},B_{1},B_{2}$ in this order, then $\mathrm{pw}(G)\leq 3$.
Proof.
As $u_{0}\notin V(P_{3})$, note that $P_{3}$ intersects the subpaths $A_{2}-u_{0},B_{1}-u_{0},B_{2}-u_{0}$. By Lemma˜4.2, $B_{1}$ cannot intersect both $A_{2}-u_{0}$ and $B_{2}-u_{0}$. By symmetry, assume that $B_{1}$ does not intersect $A_{2}-u_{0}$. Let $w$ be the vertex of $V(A_{2})\cap V(P_{3})$ which is the closest from $a_{3}$, and let $A_{3}:=P_{3}[a_{3},w]$. As $B_{1}$ does not intersect $A_{2}\setminus w$, and as $u_{0}\notin V(P_{3})$, we have $w\in V(P_{3})\cap V(A_{2})\setminus V(B_{1})$. Now, observe that if $A_{3}$ intersects the path $P_{1}$, then as $P_{3}$ is disjoint from $A_{1}$, it implies that $P_{3}$ bounces between $A_{2}$ and $B_{1}$, in which case Lemma˜4.9 applies and gives $\mathrm{pw}(G)\leq 3$. Assume now that $A_{3}$ and $P_{1}$ are disjoint. Then, we claim that we are in the configuration of Case (2), with $w$ playing the role of $u_{0}$, $P_{1}$ playing the role of $P_{3}$, and $P_{3}$ playing the role of $P_{1}$. In particular, Lemma˜4.10 then applies and gives $\mathrm{pw}(G)\leq 3$. ∎
Lemma 4.14.
If $P_{3}$ intersects the subpaths $A_{2},B_{2},B_{1}$ in this order, then $\mathrm{pw}(G)\leq 3$.
Proof.
As $u_{0}\notin V(P_{3})$, note that $P_{3}$ also intersects the subpaths $A_{2}-u_{0},B_{2}-u_{0},B_{1}-u_{0}$. We let $x,y,z$ be three distinct vertices of $P_{3}$ such that $x\in V(A_{2}),y\in V(B_{2})$ and $z\in V(A_{1})$, and such that $y$ is an internal vertex of $P_{3}[x,z]$.
Note that by Lemma˜4.2, no vertex of $P_{3}[y,b_{3}]$ belongs to $A_{2}$, and no vertex of $P_{3}[a_{3},x]$ belongs to $B_{2}$. Moreover, by Lemma˜4.9, no vertex on $P_{3}[z,b_{3}]$ belongs to $P_{2}-z$. In particular, it implies that we may assume that $x$ is the vertex from $V(A_{2})\cap V(P_{3})$ that minimizes the distance $d(a_{2},x)$, and that $y$ is the vertex from $V(B_{2})\cap V(P_{3})$ that minimizes the distance $d(b_{2},y)$.
Assume first that $B_{1}$ does not intersect $B_{2}-u_{0}$. Note that by Remark˜4.8, it implies that $y$ is in $V(P_{2})\cap V(P_{3})\setminus V(P_{1})$. We set $D_{2}:=P_{2}[y,b_{2}]$ and $D_{3}:=P_{3}[a_{3},y]$. By our minimality assumption, $y$ is must then be the only intersection of $D_{2}$ with $P_{1}\cup P_{3}$. By Lemma˜4.9, no vertex on $D_{3}$ can belong to $B_{1}$. Hence, we have that $P_{1}$ does not intersect $D_{2}$ nor $D_{3}$. By Lemma˜4.10, with $y$ playing the role of $u_{0}$, $P_{1}$ playing the role of $P_{3}$ and $D_{2}$ playing the role of $A_{1}$, we obtain that $\mathrm{pw}(G)\leq 3$.
We now assume that $B_{1}$ intersects $B_{2}-u_{0}$. In particular, $B_{1}$ does not intersect $A_{2}-u_{0}$ by Lemma˜4.2. By our choice of $x$ and Remark˜4.8, observe that $C_{2}:=P_{2}[a_{2},x]$ only intersects $P_{1}\cup P_{3}$ in $x$. We now set $A_{3}:=P_{3}[a_{3},x]$. Note that by Lemma˜4.2 and our choice of $x$, $A_{3}$ only intersects $P_{2}$ in $x$. Moreover, note that as $A_{3}$ is disjoint from $A_{1}$, if $A_{3}$ intersects $B_{1}$, then $P_{3}$ bounces between $B_{1}$ and $A_{2}$, so by Lemma˜4.9 we have $\mathrm{pw}(G)\leq 3$. We may thus assume that $x$ is the vertex in the intersection of $A_{3}$ and $P_{1}\cup P_{2}$. Note that we are then in the situation of Lemma˜4.10, with $x$ playing the role of $u_{0}$, $P_{1}$ playing the role of $P_{3}$ and $C_{2},A_{3}$ playing the respective roles of $A_{1},A_{2}$. We thus conclude that $\mathrm{pw}(G)\leq 3$. ∎
Proof of Theorem˜1.3.
As explained above, observe that, up to replacing $P_{2}$ by $P_{2}^{-1}$, one of the four cases between (1), (2), (3) and (4) should occur. We thus conclude that $\mathrm{pw}(G)\leq 3$ after applying according accordingly Lemma˜4.9, Lemma˜4.10, Lemma˜4.13 and Lemma˜4.14. ∎
5 Coverings by Isometric Subtrees
Treewidth.
A tree-decomposition of a graph $G$ is a pair $(T,\mathcal{V})$ where $T$ is a tree and $\mathcal{V}=(V_{t})_{t\in V(T)}$ is a family of subsets $V_{t}$ of $V(G)$ such that:
  • $V(G)=\bigcup_{t\in V(T)}V_{t}$;
  • for every nodes $t,t^{\prime},t^{\prime\prime}$ such that $t^{\prime}$ is on the unique path of $T$ from $t$ to $t^{\prime\prime}$, $V_{t}\cap V_{t^{\prime\prime}}\subseteq V_{t^{\prime}}$;
  • every edge $e\in E(G)$ is contained in an induced subgraph $G[V_{t}]$ for some $t\in V(T)$.
The subsets $V_{t}$ are called the bags of the tree-decomposition $(T,\mathcal{V})$, and the tree $T$ is called the decomposition tree. The width of a tree-decomposition is the maximum size of a bag minus $1$. The treewidth of a graph $G$, denoted $\mathrm{tw}(G)$ is defined as the minimum possible width of a tree-decomposition of $G$. Note that path-decompositions of $G$ correspond exactly to the tree-decompositions of $G$ whose decomposition tree is a path. In particular, we have in general $\mathrm{tw}(G)\leq\mathrm{pw}(G)$.
Recall that a graph $G$ is $2$-connected if it is connected of order at least $3$, and has no cutvertex, that is, no vertex $v$ such that GvG-v is not connected anymore. The blocks of $G$ are the inclusion-wise maximal sets of vertices $X\subseteq V(G)$ such that $G[X]$ is connected and has no cutvertex [^4] Note that here, blocks are not necessarily $2$-connected, as we allow them to have order less than $3$.. The following is a well-known and easy fact.
Proposition 5.1 (folklore).
The treewidth of a graph is equal to the maximum treewidth of its blocks (maximal $2$-connected components). Put differently, for every $v\in V(G)$, if $X_{1},\ldots,X_{k}$ denote the connected components of GvG-v, then $\mathrm{tw}(G)=\max_{1\leq i\leq k}\mathrm{tw}(G[X_{i}\cup\{v\}]).$
The following remark is immediate.
Remark 5.2.
Let $G$ be a graph edge-coverable by $k$ isometric trees, and $X$ be a block of $G$. Then $G[X]$ is also edge-coverable by $k$ isometric trees.
We now prove our main result in this section, which we restate. See 1.4
Proof.
Let $T_{1},T_{2}$ denote two isometric subtrees of $G$ that cover all its edges. By Proposition˜5.1 and Remark˜5.2, we may assume without loss of generality that $G$ is $2$-connected. In particular, $G$ has no vertex of degree $1$. Moreover, up to removing some edges incident to the leaves of $T_{1}$ and $T_{2}$, we may further assume that every edge of $T_{1}$ which is incident to a leaf of $T_{1}$ is not an edge of $T_{2}$. Let $u$ be a leaf of $T_{1}$. As $G$ has no vertex of degree $1$, $u$ must also be a vertex of $T_{2}$.
Claim 5.3.
Every edge of $G$ is vertical with respect to $u$.
Proof of the Claim.
Suppose for a contradiction that there exists some edge $vw\in E(G)$ which is not vertical with respect to $u$. Then it implies that $v$ and $w$ belong to the same layer of the BFS-layering starting from $u$, that is, $d_{G}(u,v)=d_{G}(u,w)$. We let $i\in\{1,2\}$ be such that $vw\in E(T_{i})$. As $T_{i}$ is an isometric subtree of $G$, we then also have $d_{T_{i}}(u,v)=d_{T_{i}}(u,w)$. We then obtain a contradiction as such an edge $vw$ would create a cycle in $T_{i}$. ∎
We now consider the subtree $T^{\prime}_{1}$ of $T_{1}$ formed by taking the union of all paths $P$ between $u$ and a vertex of $T_{1}$ such that all internal vertices of $P$ are in $V(T_{1})\setminus V(T_{2})$. As $G$ is $2$-connected, all leaves of $T^{\prime}_{1}$ are in $T_{2}$. Let $G_{v}$ be the subgraph of $G$ induced by the descendants of $v$, i.e., by vertices $x$ for which there exists some $ux$-path $P$ which contains $v$ and which is vertical with respect to $u$. Note that in particular, $v\in V(G_{v})$.
Claim 5.4.
Each leaf $v$ of $T^{\prime}_{1}$ which is distinct from $u$ separates $G_{v}$ from the rest of $G$.
Proof of the Claim.
Let $v$ be a leaf of $T^{\prime}_{1}$ distinct from $u$, and suppose for a contradiction, that $v$ does not separate $G_{v}$ from $G-G_{v}$. Then there exists some edge $e=xy\in E(G)$ with $x\in V(G_{v})\setminus\{v\}$ and $y\in V(G)\setminus V(G_{v})$. By ˜5.3, $e$ is vertical with respect to $u$. As we assumed that $y\notin V(G_{v})$, we must have $d_{G}(u,y)<d_{G}(u,x)$. Let $i\in\{1,2\}$ be such that $xy\in E(T_{i})$. We let $P$ denote the path in $T_{i}$ connecting $u$ to $x$, which, by ˜5.3, is vertical with respect to $u$ and thus contains $y$. As $y$ is not in $G_{v}$, observe that $P$ cannot contain $v$. See Figure˜17 for the case where $i=2$. We let $w$ be the least common ancestor in $T_{i}$ of $v$ and $x$ (recall that by definition of $T^{\prime}_{1}$, $v$ is in $V(T_{1})\cap V(T_{2})$). As $P$ avoids $v$, we have $v\neq w$. As $T_{i}$ is an isometric subgraph of $G$, we then have $d_{G}(v,x)=d_{T_{i}}(v,x)=d_{T_{i}}(v,w)+d_{T_{i}}(w,x)>d_{T_{i}}(w,x)=d_{G}(w,x)$. Now, recall that $x\in V(G_{v})$, hence there exists some $ux$-path in $G$ from $v$ to $x$ going through $v$, which is vertical with respect to $u$. In particular, as $w$ is an ancestor of $v$ in $T_{i}$, it then implies that $d_{G}(v,x)\leq d_{G}(w,x)$, giving a contradiction (see Figure˜17). ∎
Figure 17: Proof of ˜5.4 in the case where $xy\in E(T_{2})$.
As $G$ is $2$-connected, ˜5.4 implies that for each leaf $v$ of $T^{\prime}_{1}$ distinct from $u$, we have $V(G_{v})=\{v\}$. In particular, as we initially chose $u$ to be a leaf of $T_{1}$, it then implies that $T_{1}=T^{\prime}_{1}$, so the only vertices of $T_{1}$ that also belong to $T_{2}$ are its leaves. Observe that up to reproducing symmetrically the same reasoning, where the roles of $T_{1}$ and $T_{2}$ are exchanged, we may moreover assume that the only vertices of $T_{2}$ belonging to $T_{1}$ are its leaves. In other words, we are left with the case where $T_{1}$ and $T_{2}$ have no common internal nodes, and intersect exactly at their leaves.
In the remainder of the proof, we will say that a graph $H$ has a mirror-decomposition if there exist trees $T,T^{\prime}$ such that
  • $H=T\cup T^{\prime}$;
  • the trees $T$ and $T^{\prime}$ have no common internal nodes;
  • there exists a graph isomorphism $\iota$ from $T$ to $T^{\prime}$ such that for every $v\in V(T_{1})\cap V(T_{2})$, we have $\iota(v)=v$.
The proof of Theorem˜1.4 will be an immediate consequence of the following two Claims.
Claim 5.5.
$G$ has a mirror-decomposition, with respect to the trees $T_{1}$ and $T_{2}$.
Proof of the Claim.
First, observe that as $T_{1},T_{2}$ are both isometric subgraphs of $G$, for every two vertices $a,b$ in $V(T_{1})\cap V(T_{2})$, we have $d_{G}(a,b)=d_{T_{1}}(a,b)=d_{T_{2}}(a,b)$. In particular, the claim is then immediate if $T_{1}$ and $T_{2}$ have at most $2$ leaves. We now assume that $T_{1},T_{2}$ have more than $2$ leaves.
We recall that a basic property of trees is that for every three nodes $a,b,c$ in a tree $T$, there exists a unique median, that is, a unique vertex $m(a,b,c)$ of $T$ that belongs to shortest paths between each pair of $a,b$ and $c$. Note that in particular, if $a,b,c$ are pairwise distinct leaves of a tree $T$, then their median is an internal node of $T$.
In what follows, for every $i\in\{1,2\}$ and every three leaves of $T_{i}$ (and thus also of $T_{3-i}$), we let $m_{i}(a,b,c)\in V(T_{i})$ denote their median in the tree $T_{i}$. We first claim that for every $x\in\{a,b,c\}$, we have $d(m_{1}(a,b,c),x)=d(m_{2}(a,b,c),x)$. To see this, fix $a,b,c$ and for each $x\in\{a,b,c\}$, set $d_{i}(x):=d_{T_{i}}(m_{i}(a,b,c),x)$. Note first that as $T_{1},T_{2}$ are isometric subgraphs in $G$, we have for every two distinct vertices $x,y\in\{a,b,c\}$ that $d_{1}(x)+d_{1}(y)=d_{T_{1}}(x,y)=d_{T_{2}}(x,y)=d_{2}(x)+d_{2}(y)$. In particular, we then have
$2(d_{1}(a)+d_{1}(b)+d_{1}(c))$ $=d_{T_{1}}(a,b)+d_{T_{1}}(b,c)+d_{T_{1}}(a,c)$
$=d_{T_{2}}(a,b)+d_{T_{2}}(b,c)+d_{T_{2}}(a,c)$
$=2(d_{2}(a)+d_{2}(b)+d_{2}(c)).$
By a previous observation, $d_{1}(b)+d_{1}(c)=d_{2}(b)+d_{2}(c)$, hence in particular it implies that $d_{1}(a)=d_{2}(a)$. The equalities $d_{1}(b)=d_{2}(b)$ and $d_{1}(c)=d_{2}(c)$ are obtained identically.
To conclude the proof of ˜5.5, we now show that for every tree $T$, the knowledge for each triple $(a,b,c)$ of distinct leaves of the tuples
(d(m(a,b,c),a),d(m(a,b,c),b),d(m(a,b,c),c))(d(m(a,b,c),a),d(m(a,b,c),b),d(m(a,b,c),c))
can be used to reconstruct $T$ uniquely. In other words, for a fixed set $L$ of size at least $3$ and a fixed mapping $\varphi:L^{3}\to\mathbb{N}^{3}$, there exists at most one tree $T$ (up to relabelling the internal nodes) such that the leaves of $T$ are exactly the elements of $L$, and such that for every triple $(a,b,c)$ of distinct leaves, we have
$\varphi(a,b,c)=(d(m(a,b,c),a),d(m(a,b,c),b),d(m(a,b,c),c)).$
When such an equality holds for all triple of distinct leaves, we say that $T$ realizes $\varphi$. Note that by previous paragraphs proving such a statement will immediately conclude our proof of ˜5.5. We show this statement by induction on the size of $L$. If $|L|=3$, then note that the only possible tree realizing some given triple $\varphi(a,b,c)=(d_{a},d_{b},d_{c})$ is a subdivided star with three branches of size $d_{a},d_{b}$ and $d_{c}$ respectively connecting $a,b$ and $c$ to the center. Assume now that $|L|\geq 4$ and that the induction hypothesis holds. Suppose that we are given some tree $T$ with set of leaves $L$ realizing some fixed mapping $\varphi:L^{3}\to\mathbb{N}^{3}$. Let $a$ be a leaf of $T$, and let $P$ be the shortest possible path in $T$ connecting $a$ to a vertex $x$ of degree at least $3$ (such a vertex always exists as soon as $T$ has at least $3$ leaves). Then TPT-P is a tree whose set of leaves is exactly $L\setminus\{a\}$. In particular, applying induction hypothesis on $L\setminus\{a\}$ implies that TPT-P is the unique possible tree realizing the restriction of $\varphi$ on $(L\setminus\{a\})^{3}$. Moreover, as $x$ has degree at least $3$ in $T$, there exist two distinct leaves $b,c$ of TPT-P such that $x=m(a,b,c)$ in $T$. Write $\varphi(a,b,c)=(d_{a},d_{b},d_{c})$ and observe now that $x$ is necessarily the only vertex of TPT-P that lies on the $bc$-path of TPT-P such that $d(x,b)=d_{b}$ and $d(x,c)=d_{c}$. In particular, it implies that the only possibility for a tree to realize $\varphi$ is to be obtained from TPT-P after attaching some path $P^{\prime}$ of length $d_{a}$ connecting $a$ to $x$. This concludes the proof. ∎
Claim 5.6.
Every graph admitting a mirror-decomposition has treewidth at most $2$.
Proof of the Claim.
We let $H$ be a graph with a mirror-decomposition and $T,T^{\prime},\iota$ be as in the definition of mirror-decomposition. We prove by induction on $|V(T)|=|V(T^{\prime})|\geq 1$ that there exists a tree-decomposition of $H$ of width at most $2$, and such for each $v\in V(T)$, there exists a bag that contains $\{v,\iota(v)\}$. If $|V(T)|\leq 2$, then $|V(H)|\leq 3$ (unless $V(T)$ and $V(T^{\prime})$ are disjoint, in which case we also conclude immediately), and one can conclude by considering the tree-decomposition with a single bag equal to $V(H)$. Assume now that $|V(T)|=|V(T^{\prime})|\geq 3$. In particular, there exists some node $u\in V(T)$ such that TuT-u is not connected. Then, $\{u,\iota(u)\}$ forms a separator of size $2$ in $H$ (see Figure˜18).
[…]
Figure 18: Example of mirrored trees. Notice that $\{u,\iota(u)\}$ indeed forms a separator of the graph.
6 Conclusion
As mentioned in the introduction, the best lower bound we know on the possible pathwidth of graphs edge-coverable by $k$ shortest paths is $k$. Together with Theorem˜1.3, it suggests that the upper bound from Theorem˜1.2 could be potentially improved to a linear bound.
Question 6.1.
Let $G$ be a graph edge-coverable by $k$ shortest paths. Is it true that $\mathrm{pw}(G)=O(k)$? And that $\mathrm{pw}(G)\leq k$?
Another observation going in the direction of ˜6.1 is that graphs edge/vertex-coverable by $k$ shortest paths have cop number at most $k$ [AF84]. In particular, it is well-known (and not hard to prove) that the cop number of a graph is upper bounded by its treewidth, and thus also by its pathwidth.
The second question which is obviously left open by our work (and asked in [DFPT24]) concerns the existence of a polynomial upper-bound on the pathwidth of graphs which are vertex-coverable by $k$ shortest paths. At the moment, it does not seem for us that the ideas from Section˜3 generalize to this case.
As of now, we still do not know if Theorem˜1.4 generalizes in some way to larger values of $k$:
Question 6.2.
Does there exist a function $f:\mathbb{N}\to\mathbb{N}$ such that for every $k\geq 3$ and every graph $G$ edge-coverable by $k$ isometric subtrees, we have $\mathrm{tw}(G)\leq f(k)$? If yes, can $f$ be polynomial?
Note that for every $k\geq 3$, there always exists a graph edge-coverable by $k$ isometric subtrees with treewidth $k$: take any tree $T$ with at least $k$ leaves, and let $T_{1},\ldots,T_{k}$ denote $k$ disjoint copies of $T$. Let $G_{k}$ be the graph obtained after bijectively identifying for each $i\geq 2$ each leaf of $T_{i}$ with its corresponding copy in $T_{1}$. We claim that $G_{k}$ has a $K_{k,k}$-minor (contract all edges between any two internal nodes in each of the $T_{i}$’s), thus $\mathrm{tw}(G_{k})\geq k$. Moreover, this is an equality, as we claim that it is not hard to find some tree-decomposition of $G_{k}$ of width $k$ (one can reproduce the proof of ˜5.6). At the moment we do not know better lower bounds than $k$ on the possible treewidth of graphs edge-coverable by $k$ isometric subtrees.
Acknowledgements
We thank Marcin Briański for the example from Section˜1 (see Figure˜1) of graphs vertex-coverable by two isometric trees with large treewidth. We also thank Florent Foucaud for mentioning the potential algorithmic application of Theorem˜1.2 to Isometric Path Edge-Cover.